### v2

# Spectral-Operator Calculus I: Trace-Form Evaluators and Spectral Growth in the Self-Adjoint Setting

Abstract.
We develop Spectral-Operator Calculus (SOC), an axiomatic calculus for scalar evaluation of operator-generated spectral observables. This paper (SOC-I) treats the self-adjoint setting, where observables are bounded Borel transforms and locality is enforced via additivity across spectral partitions. Under unitary invariance, extensivity on orthogonal sums, projector-locality, and a dominated-convergence continuity condition, we prove a rigidity theorem on a natural trace-class envelope: every admissible evaluator agrees with a weighted trace of a single Borel nondecreasing profile applied through the functional calculus. We then introduce a spectral growth taxonomy based on eigenvalue counting asymptotics and show that the polynomial growth regime is stable under the basic constructions of the calculus. These results supply an arithmetic-neutral analytic backbone for subsequent SOC parts and for applications to concrete spectral models. A companion part treats the sectorial/holomorphic setting, where locality is formulated on log-scale via scale-band decompositions and positive kernels rather than spectral projections.
Key words and phrases:
spectral operator, functional calculus, trace-form evaluator, dominated convergence, spectral growth, admissible kernels
2020 Mathematics Subject Classification:
47A10, 47B10, 46Lxx
Series note. The SOC series develops an evaluator calculus for spectral observables generated by functional calculi. The common spine is an axiomatics for scalar evaluators (unitary invariance, extensivity, localization under coarse-graining, dominated-continuity, and normalization), together with rigidity and stability consequences that are designed to persist under admissible limiting schemes.
SOC-I develops the self-adjoint/PVM-localized model case. It fixes the evaluator axioms, proves a trace-form rigidity theorem on a natural trace-class envelope, and introduces a spectral growth taxonomy and continuity framework for scalar evaluations built from spectra. A companion part develops the sectorial/holomorphic model case (Haase-type functional calculus; see Haase [1]), where the localization gate is implemented on log-scale by scale-band decompositions and kernel positivity rather than by spectral projections.
Later parts build on these model cases to study admissible kernels and cutoff schemes, analytic continuation procedures, and boundary and reconstruction questions.
1. Introduction
The basic objects in this paper are self-adjoint operators, their spectral measures, and bounded Borel transforms. Concretely, we work on a fixed complex separable Hilbert space $H$, regard self-adjoint operators $D$ on $H$ as models, and treat bounded Borel functions $f:\mathbb{R}\to\mathbb{C}$ as generating bounded operators $f(D)$ through the functional calculus provided by the spectral theorem. These operators $f(D)$ represent the spectral geometries on which our scalar-valued functionals act. In this sense, all primitive data in the calculus are operators and spectra, not bare scalar sequences.
At the level of structure, SOC is organized around three interacting gates: (i) a localization rule that specifies how evaluator values decompose under coarse-graining; (ii) a dominated-continuity rule that legalizes passage to limits inside the evaluator; and (iii) a normalization/nontriviality rule that excludes degenerate assignments. The operator class and functional calculus supply the observable family and the available coarse-grainings. In the self-adjoint setting of this part, coarse-graining is spectral: we partition the spectrum and demand additivity across the corresponding spectral projections. In the sectorial setting, spectral projections are unavailable in general; SOC instead uses scale-band decompositions $t\mapsto\psi(tA)$ and controls interference through a positive kernel on log-scale. This paper develops the PVM-localized model case and proves that the gates force trace-form rigidity on the trace-class envelope.
1.1. Main results and contributions
The central structural object of the paper is an evaluator
$\mathcal{E}:\mathcal{C}\to[0,\infty]$
defined on a class $\mathcal{C}$ of bounded spectral geometries. The class $\mathcal{C}$ is required to be stable under bounded Borel transforms and under the operations needed later for examples and continuity statements. The evaluator $\mathcal{E}$ is required to satisfy natural axioms: unitary invariance, locality on spectral projections, extensivity on orthogonal sums, a dominated-convergence continuity property, and a mild growth condition that prevents trivialization (see Axiom A5 in Section 3.1). The guiding question is whether such an abstract evaluator is forced to be of trace form, and, if so, how rigid that representation is.
Our first main result is Theorem 3.1, a trace-form representation theorem for evaluators. Under the standing hypotheses on $\mathcal{C}$ and the evaluator axioms, there exist a Borel-measurable nondecreasing function $h:\mathbb{R}\to\mathbb{R}$ and a normalization constant $c_{\mathcal{E}}\geq 0$ such that $\mathcal{E}(f(D))$ is given by the trace of $h(f(D))$ whenever this operator is trace class. In particular, if we restrict $\mathcal{E}$ to the trace-class envelope
$\mathcal{C}_{1}:=\mathcal{C}\cap\mathcal{S}_{1}(H),$
then every admissible value $\mathcal{E}(X)$ with $X\in\mathcal{C}_{1}$ is completely determined by the spectral geometry of $X$ together with the fixed profile $h$ and the overall scale $c_{\mathcal{E}}$. In this sense the evaluator has no additional degrees of freedom once the profile and normalization are fixed.
The second main result is a spectral growth taxonomy for self-adjoint operators. We work with counting functions $N_{D}(\Lambda)$ and related spectral distributions, and use their asymptotic behaviour as $\Lambda\to\infty$ to organize operators into growth classes. In particular, we single out a polynomial growth regime that captures the standard elliptic and Laplace-type models and show in Theorem 3.5 that this class is stable under the basic constructions of the calculus, including monotone regularly varying functional calculus, finite orthogonal sums, and tensor products. This growth framework is then used in examples and in the continuity theory for evaluators in Section 5.
A further theme of the paper is continuity under natural limiting procedures. We study approximations of spectral geometries by truncations and cutoffs and show that the evaluator values behave well under dominated limits. In particular, we prove that monotone and dominated approximations by bounded transforms do not introduce spurious boundary contributions in the trace-form representation.
Conceptually, the contribution of this paper is twofold. First, at the level of scalar evaluation, the axioms force a rigidity phenomenon on the trace-class envelope: any evaluator satisfying the standing structural conditions must be given by a single nondecreasing profile applied through the functional calculus and traced, so that scalar values carry no independent structure beyond what is encoded in the spectra and the chosen profile. Second, on the geometric side, the spectral growth classification and closure results provide a stable regime in which the spectral calculus and the trace-form representation interact well under the operations needed in later parts of the series.
Throughout, the presentation is self-contained and arithmetic-neutral. All statements are made at the level of operators, spectra, and traces, without appealing to arithmetic or number-theoretic input. The intent is that the evaluator representation, growth taxonomy, and continuity framework developed here can be reused across different spectral models and applications without revisiting these analytic foundations.
1.2. Relation to prior work
There is a substantial literature on trace ideals, unitarily invariant norms, and operator inequalities, with classical references including Reed-Simon [2], Simon’s monograph on trace ideals [3], and Bhatia’s text on matrix analysis [4]. In that setting one typically starts from a given trace and studies the properties of associated norms and functionals on operator ideals. The perspective adopted here is different: we start from an abstract evaluator $\mathcal{E}$ defined on a class $\mathcal{C}$ of spectral geometries, subject only to unitary invariance, projector-locality, dominated continuity, and a growth bound, and obtain a universal trace-form representation for such evaluators on a trace-class envelope. In this sense Theorem 3.1 can be viewed as a representation theorem for evaluator functionals, rather than as a study of a fixed underlying trace.
Classical uniqueness theorems for symmetric gauge functionals work in the opposite direction: they assume the trace or norm structure is given a priori and then classify functionals compatible with it, whereas here trace form itself is derived as a consequence of the evaluator axioms, without presupposing any trace or ideal structure.
On the functional-calculus side, SOC is modular: it takes as input an operator class equipped with a functional calculus and a compatible notion of coarse-graining, and it studies which scalar evaluators are forced by invariance, locality, and dominated-limit legality. SOC-I instantiates this program in the self-adjoint/Borel setting, where the spectral theorem supplies both the calculus and the canonical coarse-grainings (spectral partitions via the PVM). Haase’s framework for sectorial operators [1] supplies the corresponding input for a second SOC model case based on holomorphic functional calculus, where PVM locality is replaced by scale-band decompositions and kernel positivity on log-scale. The novelty in SOC is not the existence of these calculi, but the rigidity and stability constraints on evaluators that follow once a localization structure is fixed.
On the growth side, the taxonomy introduced here is influenced by the language of regular variation and spectral asymptotics. Counting-function techniques and regularly varying functions are standard tools in the analysis of eigenvalue distributions, as in the work of Bingham-Goldie-Teugels [5] and in more recent treatments of spectral densities and hypertraces in noncommutative geometry [6], as well as in explicit counting-function studies for fractal sprays and related models [7]. Our contribution is to formulate a growth classification directly in terms of counting functions and to prove stability of the polynomial class under the structural operations required by the evaluator calculus, such as functional calculus, orthogonal sums, and tensor products, without committing to a specific geometric model.
This paper should therefore be read as an operator-theoretic calculus for evaluators and spectral growth, rather than as a contribution to the fine structure of trace ideals or spectral asymptotics. It does not aim to optimize asymptotic constants or to exhaust the classical theory; instead, it isolates a structural setting in which evaluators, spectral geometries, and growth conditions fit together coherently, and in which the main representation and continuity results can be applied across different models without further analytic reconstruction. The results of SOC-I are intended to be reusable: they provide an analytic backbone for later work on kernels, cutoff schemes, and flow constructions on spectral data, while the present part can be read independently of any later developments in the series.
1.3. Organization of the paper
Section 2 collects the spectral calculus, evaluator notation, and the standing assumptions on the stable class $\mathcal{C}$. Section 3 states the evaluator axioms and the two main results: the trace-form representation theorem and the spectral growth taxonomy. Section 4 proves these results, delegating measure-theoretic details to Appendix A and counting-function estimates to Appendix B. Section 5 illustrates the continuity theory for evaluators and gives examples of spectral geometries in the growth classes introduced in Section 3.
2. Preliminaries
Throughout, $H$ denotes a complex separable Hilbert space with inner product $\langle\cdot,\cdot\rangle$ linear in the second argument and norm $\|x\|=\sqrt{\langle x,x\rangle}$. All operators are densely defined and self-adjoint unless explicitly stated otherwise. For a self-adjoint operator $D$ on $H$, we write $\sigma(D)$ for its spectrum and $E_{D}$ for the associated projection-valued spectral measure on the Borel subsets of $\mathbb{R}$.
2.1. Spectral measures and functional calculus
Let $D$ be self-adjoint on $H$. By the spectral theorem, there exists a unique projection-valued measure $E_{D}$ on the Borel $\sigma$-algebra of $\mathbb{R}$ such that
$D=\int_{\mathbb{R}}\lambda\,dE_{D}(\lambda)$ (2.1)
in the sense of strong operator convergence on the domain of $D$; see Reed-Simon [2]*{Theorem VIII.3}. For any bounded Borel function $f:\mathbb{R}\to\mathbb{C}$, the bounded operator
$f(D):=\int_{\mathbb{R}}f(\lambda)\,dE_{D}(\lambda)$ (2.2)
is defined by spectral calculus. We will refer to such operators $f(D)$, with $f$ bounded Borel, as spectral geometries associated with $D$. In particular, the basic objects of the calculus will be self-adjoint operators $D$, their spectral measures $E_{D}$, and the bounded spectral geometries $f(D)$ built from them. If $f$ is real-valued, then $f(D)$ is self-adjoint. The assignment $f\mapsto f(D)$ is a unital *-homomorphism from the algebra of bounded Borel functions on $\mathbb{R}$ into the bounded operators on $H$.
For vectors $x,y\in H$, the scalar measure
$\mu_{x,y}(B):=\langle E_{D}(B)x,y\rangle,\qquad B\subset\mathbb{R}\ \text{Borel},$
is complex-valued and countably additive; when $x=y$ we write $\mu_{x}:=\mu_{x,x}$, a finite positive measure. For bounded Borel $f$,
$\langle f(D)x,y\rangle=\int_{\mathbb{R}}f(\lambda)\,d\mu_{x,y}(\lambda).$
When $D$ has pure-point spectrum with finite multiplicities, we write $\{\lambda_{n}(D)\}_{n\geq 1}$ for its eigenvalues (repeated according to multiplicity) and choose an orthonormal basis of eigenvectors. Then $D$ is unitarily equivalent to a diagonal operator on $\ell^{2}$, and for bounded Borel $f$,
$f(D)\sim\operatorname{diag}\big(f(\lambda_{1}(D)),f(\lambda_{2}(D)),\ldots\big).$
Remark 2.1 (Borel functional calculus).
In what follows, whenever a Borel-measurable function $h:\mathbb{R}\to\mathbb{R}$ is applied to a self-adjoint operator $D$, the operator $h(D)$ is understood in the sense of the Borel functional calculus described above. In particular, Theorem 3.1 will invoke $h(f(D))$ with $f$ bounded Borel and $h$ Borel-measurable and nondecreasing; no continuity of $h$ beyond Borel measurability is required to define $h(f(D))$.
2.2. Trace class operators and traces
Let $\mathcal{S}_{1}(H)$ denote the ideal of trace-class operators on $H$. For $T\in\mathcal{S}_{1}(H)$, its trace is defined by
$\operatorname{Tr}(T)=\sum_{n\geq 1}\langle Te_{n},e_{n}\rangle,$
where $\{e_{n}\}$ is any orthonormal basis of $H$; the value is independent of the basis. If $T$ is positive trace class, $\operatorname{Tr}(T)$ coincides with the sum of its eigenvalues counted with multiplicity.
For a self-adjoint $D$ and bounded Borel $f$, the operator $f(D)$ need not be trace class. Whenever $f(D)\in\mathcal{S}_{1}(H)$, the spectral theorem gives
$\operatorname{Tr}(f(D))=\int_{\mathbb{R}}f(\lambda)\,d\nu_{D}(\lambda)$ (2.3)
where the spectral trace measure $\nu_{D}$ is defined by
$\nu_{D}(B)=\operatorname{Tr}(E_{D}(B)),\qquad B\subset\mathbb{R}\ \text{Borel},$
whenever this expression is finite. In many situations, $\nu_{D}$ is infinite on unbounded sets but finite on compact sets; in that case the above formula is interpreted under explicit integrability or cutoff hypotheses stated later.
A regularized trace on an operator $A$ is any assignment $A\mapsto\operatorname{Tr}_{\mathrm{reg}}(A)$ defined on a class of operators that extends $\operatorname{Tr}$ on $\mathcal{S}_{1}(H)$ and is compatible with spectral calculus. In this paper we only use spectral-cutoff regularization. For each self-adjoint $D$ we fix a family of bounded Borel functions $\chi_{\Lambda}:\mathbb{R}\to[0,1]$ with $\chi_{\Lambda}(\lambda)=1$ for $|\lambda|\leq\Lambda$, $\chi_{\Lambda}(\lambda)=0$ for $|\lambda|\geq 2\Lambda$, and $\chi_{\Lambda}(\lambda)\to 1$ pointwise as $\Lambda\to\infty$. Let
$T_{\Lambda}(D)=\chi_{\Lambda}(D)$
be the corresponding spectral cutoff operators. Whenever the limit exists, we define the spectral-cutoff trace of $f(D)$ by
$\operatorname{Tr}_{\mathrm{cut}}(f(D))=\lim_{\Lambda\to\infty}\operatorname{Tr}\big(T_{\Lambda}(D)\,f(D)\big).$
Whenever we speak of a regularized trace or a compatible regularized trace in the sequel, we mean this spectral-cutoff trace $\operatorname{Tr}_{\mathrm{cut}}$, taken with respect to the fixed family $\{T_{\Lambda}(D)\}$. No other regularization schemes are used in this paper.
2.3. Counting functions and growth
When $D$ has discrete spectrum with finite multiplicities accumulating only at $\pm\infty$, its eigenvalues can be numbered as $\{\lambda_{n}(D)\}_{n\geq 1}$, and the associated counting function is
$N_{D}(\Lambda):=\#\{n\geq 1:|\lambda_{n}(D)|\leq\Lambda\},\qquad\Lambda\geq 0,$
where each eigenvalue is counted with multiplicity.
Convention 2.2 (Counting-function regularity).
Whenever we work with a counting function $N_{D}$, we assume that $D$ has discrete spectrum with finite multiplicities, so that
$N_{D}(\Lambda)<\infty\qquad\text{for all }\Lambda\geq 0,$
and that $N_{D}(\Lambda)$ is eventually nondecreasing as $\Lambda\to\infty$; that is, there exists $\Lambda_{0}\geq 0$ such that
$\Lambda_{1}\leq\Lambda_{2},\ \Lambda_{1},\Lambda_{2}\geq\Lambda_{0}\quad\Longrightarrow\quad N_{D}(\Lambda_{1})\leq N_{D}(\Lambda_{2}).$
The asymptotic behavior of $N_{D}(\Lambda)$ as $\Lambda\to\infty$ defines the growth class of $D$. Typical examples include polynomial growth
$N_{D}(\Lambda)\sim C\,\Lambda^{d},\qquad\Lambda\to\infty,$
stretched-exponential behavior
$N_{D}(\Lambda)\sim\exp\big(c\,\Lambda^{\alpha}\big),\qquad c>0,\ 0<\alpha<1,$
and logarithmic or slower growth. These classes will be formalized in Section 3 and shown to be stable under the structural operations needed for the spectral calculus.
Remark 2.3 (Counting functions and spectral measures).
Three related objects will appear repeatedly:
- The counting function $N_{D}(\Lambda)$ counts eigenvalues of a discrete-spectrum operator $D$ up to a cutoff $\Lambda$. - The trace spectral measure $\nu_{D}$ of $D$ is defined (when finite) by
$\nu_{D}(B):=\operatorname{Tr}(E_{D}(B)),\qquad B\subset\mathbb{R}\ \text{Borel},$
and encodes multiplicities via the spectral projections $E_{D}(B)$. - Given an evaluator $\mathcal{E}$, the evaluator measure $\mu_{D}$ is defined (when $E_{D}(B)\in\mathcal{C}$) by
$\mu_{D}(B):=\mathcal{E}(E_{D}(B)).$
In the growth taxonomy of Section 3.3 and Appendix B we work primarily with $N_{D}$. In Appendix A and Section 4.1 we work with the scalar measures $\nu_{D}$ and $\mu_{D}$ attached to the spectral measure $E_{D}$; there we use the Radon-Nikodym theorem to express $\mu_{D}$ as a density with respect to $\nu_{D}$ (see Remark A.8). These three objects carry complementary information and will not be interchanged: $N_{D}$ is a counting function, $\nu_{D}$ encodes the ordinary trace, and $\mu_{D}$ encodes the evaluator.
2.4. Spectral preorder and majorization
Let $X$ and $Y$ be self-adjoint operators on $H$ with discrete spectra and finite multiplicities. Write $\{\lambda_{n}(X)\}_{n\geq 1}$ and $\{\lambda_{n}(Y)\}_{n\geq 1}$ for their eigenvalues, counted with multiplicity and arranged in nonincreasing order of absolute value:
$|\lambda_{1}(X)|\geq|\lambda_{2}(X)|\geq\cdots,\qquad|\lambda_{1}(Y)|\geq|\lambda_{2}(Y)|\geq\cdots.$
We write
$\lambda_{n}^{\downarrow}(X):=\text{the $n$-th eigenvalue of $X$ in this ordering}.$
Definition 2.4 (Spectral preorder).
For self-adjoint $X,Y$ with discrete spectra as above, we write
$X\preceq Y$
if for every $k\geq 1$,
$\sum_{n=1}^{k}\lambda_{n}^{\downarrow}(X)\;\leq\;\sum_{n=1}^{k}\lambda_{n}^{\downarrow}(Y).$
This preorder is invariant under unitary conjugation and coincides with the usual notion of spectral majorization; see, for example, Fan [8] and Kosaki [9] for background and equivalent formulations (e.g. in terms of Ky Fan norms and convex trace functionals).
When $X$ and $Y$ are positive and trace class, the spectral preorder $\preceq$ interacts naturally with unitarily invariant trace functionals. In particular, if $X\preceq Y$ and $\Phi:[0,\infty)\to[0,\infty)$ is convex and nondecreasing with $\Phi(0)=0$, then
$\operatorname{Tr}\big(\Phi(X)\big)\;\leq\;\operatorname{Tr}\big(\Phi(Y)\big),$
whenever both traces are finite. This type of inequality underlies the spectral monotonicity properties used later in the growth taxonomy and evaluator arguments.
Remark 2.5 (Scope of the preorder in this paper).
In this paper we use the spectral preorder $\preceq$ only for operators with discrete spectra and finite multiplicities, as in Definition 2.4. All subsequent appearances of $\preceq$ refer to that setting. Extensions of spectral majorization to more general self-adjoint operators via strong resolvent limits are well developed in the literature (see, for example, Bhatia [4]*Chapter III), but we do not rely on any such extension here, and no properties of $\preceq$ beyond the discrete-spectrum case are needed for the results of Section 3, Section 4, and Section 5.
2.5. Evaluators and structural notation
A spectral geometry (or spectral transform) of a self-adjoint operator $D$ is any bounded Borel transform $f(D)$ obtained from $D$ via the functional calculus of Section 2.1. Thus
$f(D)=\int_{\mathbb{R}}f(\lambda)\,dE_{D}(\lambda)$
for some bounded Borel function $f:\mathbb{R}\to\mathbb{C}$, and $f(D)$ is a bounded operator on $H$. When $f$ is real-valued, $f(D)$ is self-adjoint.
We will work with a fixed stable class $\mathcal{C}$ of bounded spectral geometries. By this we mean a family $\mathcal{C}$ of self-adjoint bounded operators on $H$ satisfying:
  1. 1.
    If $X\in\mathcal{C}$ and $f:\mathbb{R}\to\mathbb{R}$ is bounded and Borel measurable, then $f(X)\in\mathcal{C}$. (Since $f$ is real-valued and $X$ is self-adjoint, the spectral calculus of Section 2.1 ensures that $f(X)$ is again self-adjoint.)
  2. 2.
    If $X,Y\in\mathcal{C}$ act on orthogonal subspaces and $X\oplus Y$ denotes their block-diagonal sum, then $X\oplus Y\in\mathcal{C}$.
  3. 3.
    If $X_{n}\in\mathcal{C}$ is a monotone sequence of self-adjoint operators converging strongly to $X$, then $X\in\mathcal{C}$.
In particular, $\mathcal{C}$ is closed under bounded Borel functional calculus, finite orthogonal sums, and strong limits of monotone sequences whenever the limit exists. These closure properties will be part of the standing assumptions on the stable class throughout the paper.
An evaluator on $\mathcal{C}$ is a map
$\mathcal{E}:\mathcal{C}\to[0,\infty]$
assigning a nonnegative extended real number to each bounded spectral geometry in $\mathcal{C}$. The structural axioms imposed on evaluators - unitary invariance, extensivity on orthogonal sums, projector-locality, dominated continuity, and a normalization/growth bound - will be stated and fixed in Section 3.1 as Axioms A1-A5. Until then, $\mathcal{E}$ is only used as notation for a generic scalar-valued functional on $\mathcal{C}$.
These conventions complete the preparatory framework for the evaluator calculus: Section 3 will formalize the axioms satisfied by $\mathcal{E}$ and state the main representation and growth results under those assumptions.
3. Evaluator axioms and main results
In this section we fix the structural hypotheses on the evaluator and on the class $\mathcal{C}$ of spectral geometries on which it acts, and we state the two main theorems of the paper. The first result is a trace-form representation theorem: any evaluator satisfying the axioms below is of trace form on a natural trace-class envelope $\mathcal{C}_{1}$. The second result is a spectral growth taxonomy for self-adjoint operators, together with basic closure properties of the corresponding growth classes.
3.1. Standing assumptions and axioms
In this section we fix the structural framework for the evaluator calculus used throughout the paper. All later results in SOC-I are formulated under these standing assumptions.
We work on a fixed complex separable Hilbert space $H$. All operators are understood to act on $H$, and all spectral measures and functional calculi are taken with respect to this space. For a self-adjoint operator $D$ with spectral measure $E_{D}(\cdot)$ and any bounded Borel function $f:\mathbb{R}\to\mathbb{R}$, the bounded transform
$f(D)=\int_{\mathbb{R}}f(\lambda)\,dE_{D}(\lambda)$
is defined via the spectral calculus of Section 2.1.
We consider a class $\mathcal{C}$ of bounded spectral geometries built from self-adjoint operators and bounded Borel functions, in the following sense.
  • A typical element of $\mathcal{C}$ has the form $f(D)$, where $D$ is a self-adjoint operator on $H$ and $f:\mathbb{R}\to\mathbb{R}$ is bounded and Borel measurable, with
    $f(D)=\int_{\mathbb{R}}f(\lambda)\,dE_{D}(\lambda)$
    defined via the spectral calculus of Section 2.1.
  • In particular, every $X\in\mathcal{C}$ is a bounded self-adjoint operator on $H$.
  • The class $\mathcal{C}$ is assumed to be stable under the operations needed in the sequel: if $X\in\mathcal{C}$ and $g:\mathbb{R}\to\mathbb{R}$ is bounded Borel, then $g(X)\in\mathcal{C}$; if $X,Y\in\mathcal{C}$ act on orthogonal subspaces, then $X\oplus Y\in\mathcal{C}$; and if $X_{n}\in\mathcal{C}$ is a monotone sequence converging strongly to $X$, then $X\in\mathcal{C}$.
We write
$\mathcal{C}_{1}:=\mathcal{C}\cap\mathcal{S}_{1}(H)$
for the trace-class envelope of $\mathcal{C}$, where $\mathcal{S}_{1}(H)$ denotes the trace-class operators on $H$. The trace-form representation theorem will show that on $\mathcal{C}_{1}$ any evaluator satisfying the axioms below must be of trace form. Outside $\mathcal{C}_{1}$ we do not assert that $\mathcal{E}$ has trace form.
We equip $\mathcal{C}$ with an evaluator
$\mathcal{E}:\mathcal{C}\to[0,\infty]$
satisfying the following axioms.
Axiom A1 (Unitary invariance).
If $U$ is unitary and $X\in\mathcal{C}$, then
$\mathcal{E}(U^{*}XU)=\mathcal{E}(X).$
Axiom A2 (Extensivity on orthogonal sums).
If $X,Y\in\mathcal{C}$ act on orthogonal subspaces and if $X\oplus Y\in\mathcal{C}$ denotes their block-diagonal sum, then
$\mathcal{E}(X\oplus Y)=\mathcal{E}(X)+\mathcal{E}(Y).$
Axiom A3 (Projector locality).
If $X\in\mathcal{C}$ has a finite spectral decomposition
$X=\sum_{j=1}^{m}\lambda_{j}P_{j}$
with pairwise orthogonal spectral projections $P_{j}$ and real scalars $\lambda_{j}$, then
$\mathcal{E}(X)=\sum_{j=1}^{m}\mathcal{E}(\lambda_{j}P_{j}),$
whenever the right-hand side makes sense in $0,\infty]$. In particular, the value of $X$ is determined by its values on spectral projector blocks.
Axiom A4 (Dominated continuity in the strong operator topology).
Let $X_{n},X,Y\in\mathcal{C}$ be positive operators such that
  • $X_{n}\to X$ strongly, and
  • $0\leq X_{n}\leq Y$ for all $n$,
with $\mathcal{E}(Y)<\infty$. Then
$\mathcal{E}(X_{n})\to\mathcal{E}(X).$
Here $X_{n}\to X$ strongly means convergence in the strong operator topology. In applications we will use Axiom A4 primarily for bounded spectral geometries $f(D)$ with $f$ bounded Borel, where it justifies the monotone limit and dominated-convergence arguments in the proof of the trace-form representation and in the continuity examples of Section 5.
Axiom A5 (Normalization and growth bound).
We assume that the identity operator $I$ belongs to $\mathcal{C}$, that
$0<\mathcal{E}(I)<\infty,$
and that for every scalar $t\geq 0$ the scalar multiple $tI$ lies in $\mathcal{C}$ and is evaluated by
$\mathcal{E}(tI)=t\,\mathcal{E}(I).$
Thus the evaluator is nontrivial and finite on the positive scalar multiples of the identity, and these operators provide a harmless growth bound for dominated-convergence arguments (for example, when comparing a positive $X\in\mathcal{C}$ with $Y=\|X\|_{\mathrm{op}}\,I$).
These axioms fix the structural setting for the trace-form representation theorem and the growth taxonomy developed in the rest of the paper. In particular, they encode the requirements that any lawful evaluator must satisfy in order to behave well under unitary conjugation, orthogonal decomposition, spectral partitions, and dominated limits. In terms of the three-gate structure of Section 1: Axioms A1 and A3 implement the localization gate; Axiom A4 implements dominated-continuity; and Axiom A5 implements normalization, while Axiom A2 (extensivity) is a structural compatibility condition. Under Axioms A1-A5, Theorem 3.1 will show that on $\mathcal{C}_{1}$ the evaluator $\mathcal{E}$ is necessarily of trace form with a single Borel nondecreasing profile and a global normalization constant, and Section 3.3 will introduce the growth classes compatible with this framework.
3.2. Trace-form representation of evaluators
We now state the first main theorem. Under the standing assumptions of Section 3.1, any evaluator $\mathcal{E}$ on a stable class $\mathcal{C}$ of spectral geometries that satisfies Axioms A1-A5 is necessarily of trace form on the trace-class envelope $\mathcal{C}_{1}$.
Theorem 3.1 (Trace-form representation).
Let $\mathcal{C}$ and $\mathcal{E}$ satisfy the standing assumptions of Section 3.1, and write
$\mathcal{C}_{1}:=\mathcal{C}\cap\mathcal{S}_{1}(H)$
for the trace-class envelope of $\mathcal{C}$. Then there exist
- a Borel-measurable, nondecreasing function $h:\mathbb{R}\to\mathbb{R}$ with $h(0)=0$, and - a constant $c_{\mathcal{E}}>0$,
such that the following holds.
1. For every positive operator $X\in\mathcal{C}_{1}$,
$\mathcal{E}(X)=c_{\mathcal{E}}\;\operatorname{Tr}\big(h(X)\big).$
2. More generally, for every self-adjoint operator $D$ on $H$ and every bounded Borel function $f:\mathbb{R}\to\mathbb{R}$ such that $f(D)\in\mathcal{C}_{1}$,
$\mathcal{E}\big(f(D)\big)=c_{\mathcal{E}}\;\operatorname{Tr}\big(h(f(D))\big).$
Here $h(X)$ and $h(f(D))$ are defined by the Borel functional calculus of Section 2.1, and the traces are finite for all $X\in\mathcal{C}_{1}$ by construction of $\mathcal{C}_{1}$.
Moreover, if $\tilde{h}$$\mathbb{R}\to\mathbb{R}$ is another Borel-measurable, nondecreasing function and $\tilde{c}_{\mathcal{E}}>0$ is a constant such that
$\mathcal{E}(X)=\tilde{c}_{\mathcal{E}}\;\operatorname{Tr}\big(\tilde{h}(X)\big)\qquad\text{for all }X\in\mathcal{C}_{1},$
then there exist $a>0$ and $b\in\mathbb{R}$ such that
$\tilde{h}=a\,h+b\quad\text{and}\quad\tilde{c}_{\mathcal{E}}=c_{\mathcal{E}}/a.$
In particular, on $\mathcal{C}_{1}$ the pair $(h,c_{\mathcal{E}})$ is unique up to multiplication of $h$ by a positive scalar and addition of a constant.
Proof.
See Section 4.1 and Appendix A. ∎
Remark 3.2 (Scope of Theorem 3.1).
Theorem 3.1 applies only to the trace-class envelope $\mathcal{C}_{1}:=\mathcal{C}\cap\mathcal{S}_{1}(H)$ of the evaluator domain. In particular, if $X\in\mathcal{C}$ is not trace class (for example, the identity on an infinite-dimensional space or a diagonal operator with slowly decaying eigenvalues), the theorem does not assert that $\mathcal{E}(X)$ can be written as a trace $\operatorname{Tr}(h(X))$ for any profile $h$. Non-trace-class spectral geometries enter the present part only through the growth taxonomy of Section 3.3 and the continuity theory of Section 4; any extension of evaluators to them via regularization lies outside the scope of this paper. The dominated-continuity hypothesis A4 is tailored to this setting and will be used in Section 4 and Section 5.5 for bounded spectral geometries $f(D)$ with $f$ bounded Borel, where it justifies the monotone limit and dominated-convergence arguments in the proof of the representation theorem and in the continuity examples.
Remark 3.3 (Evaluator rigidity on $\mathcal{C}_{1}$).
Theorem 3.1 shows that, on the trace-class envelope $\mathcal{C}_{1}$, the evaluator $\mathcal{E}$ has no additional degrees of freedom beyond the choice of the profile $h$ and the normalization constant $c_{\mathcal{E}}$. Once $\mathcal{C}$, the axioms A1-A5, and the overall scale are fixed, every admissible value $\mathcal{E}(X)$ with $X\in\mathcal{C}_{1}$ is given by $\operatorname{Tr}(h(X))$ up to the constant factor $c_{\mathcal{E}}$. In particular, any other evaluator $\tilde{\mathcal{E}}$ satisfying the same axioms and finite on $\mathcal{C}_{1}$ differs only by a change of normalization.
3.3. Spectral growth classes and closure
We next formulate the spectral growth taxonomy. Throughout this subsection, $D$ denotes a self-adjoint operator on $H$ with discrete spectrum and counting function
$N_{D}(\Lambda):=\#\{n\geq 1:|\lambda_{n}(D)|\leq\Lambda\},\qquad\Lambda\geq 0,$
where eigenvalues are counted with multiplicity. The asymptotic behaviour of $N_{D}(\Lambda)$ encodes the large-scale structure of the spectral geometry generated by $D$; we use it to organize operators (and their associated spectral geometries $f(D)$ with bounded Borel $f$) into growth classes.
Definition 3.4 (Growth classes of spectral geometries).
Let $D$ be a self-adjoint operator with discrete spectrum and counting function $N_{D}(\Lambda)$.
1. $D$ has polynomial growth (of order $d>0$) if
$N_{D}(\Lambda)\sim C\,\Lambda^{d}\quad\text{as }\Lambda\to\infty$ (3.1)
for some constant $C>0$.
2. $D$ has stretched-exponential growth if there exist constants $c>0$ and $0<\alpha<1$ such that
$N_{D}(\Lambda)\sim\exp\big(c\,\Lambda^{\alpha}\big)\quad\text{as }\Lambda\to\infty.$ (3.2)
3. $D$ has logarithmic (or slower) growth if $N_{D}(\Lambda)\to\infty$ and either
$N_{D}(\Lambda)\sim c\,\log\Lambda\quad\text{as }\Lambda\to\infty$ (3.3)
for some constant $c>0$, or
$N_{D}(\Lambda)=o(\log\Lambda)\quad\text{as }\Lambda\to\infty.$
We write $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\mathrm{log}}$ for the families of self-adjoint operators (and the spectral geometries they generate) with polynomial, stretched-exponential, and logarithmic-or-slower growth, respectively. These families are disjoint: no operator can belong to more than one class; see Section 4.2 and Appendix B for a justification based on regular variation.
Theorem 3.5 (Closure of spectral growth classes).
Let $D,E$ be self-adjoint operators with discrete spectra and counting functions $N_{D},N_{E}$.
1. Monotone functional calculus (polynomial case). Let $f:[0,\infty)\to[0,\infty)$ be monotone increasing, unbounded, and regularly varying of positive index. If $D\in\mathcal{G}_{\mathrm{poly}}$, then the spectral transform $f(D)$ also lies in $\mathcal{G}_{\mathrm{poly}}$.
2. Finite orthogonal sums. If $D,E$ lie in the same growth class, then $D\oplus E$ lies in that class, and
$N_{D\oplus E}(\Lambda)=N_{D}(\Lambda)+N_{E}(\Lambda),\qquad\Lambda\geq 0.$
3. Tensor products (polynomial case). Suppose $D,E\in\mathcal{G}_{\mathrm{poly}}$ and for every $\varepsilon>0$ there exist constants $C_{D,\varepsilon},C_{E,\varepsilon}>0$ such that
$N_{D}(\Lambda)\leq C_{D,\varepsilon}\,\Lambda^{d_{D}+\varepsilon},\qquad N_{E}(\Lambda)\leq C_{E,\varepsilon}\,\Lambda^{d_{E}+\varepsilon}$
for all sufficiently large $\Lambda$. Then for every $\varepsilon>0$ there exists $C_{\varepsilon}>0$ such that
$N_{D\otimes E}(\Lambda)\leq C_{\varepsilon}\,\Lambda^{d_{D}+d_{E}+\varepsilon}$
for all sufficiently large $\Lambda$. In particular, if $D,E\in\mathcal{G}_{\mathrm{poly}}$ then $D\otimes E\in\mathcal{G}_{\mathrm{poly}}$.
Proof.
See Section 4.2 and Appendix B. ∎
Remark 3.6 (Nonpolynomial growth regimes).
The monotone functional-calculus and tensor-product arguments in Section 4.2 extend, under mild additional regularity assumptions on the asymptotics of $N_{D}$, to stretched-exponential and logarithmic-or-slower growth. In particular, Lemma B.5 and the regular-variation calculus show that if $N_{D}$ has stretched-exponential (respectively logarithmic-or-slower) growth and $f$ is monotone and regularly varying, then the transformed counting function $N_{f(D)}$ has the same qualitative asymptotic type. Since these regimes are not used elsewhere in this part, we do not spell out full statements or proofs; see Appendix B and Bingham-Goldie-Teugels [5] for background on regularly varying functions.
Theorem 3.5 shows that, in particular, the polynomial growth class $\mathcal{G}_{\mathrm{poly}}$ is stable under the basic constructions of the spectral calculus that we use later. This stability under monotone functional calculus, orthogonal sums, and tensor products underlies the continuity and evaluator examples in Section 4 and Section 5.
4. Proofs of the main results
In this section we prove the representation and growth theorems stated in Section 3. Section 4.1 outlines the proof of Theorem 3.1, with the main construction deferred to Appendix A. Section 4.2 proves the closure of the growth families in Theorem 3.5, with supporting estimates in Appendix B.
4.1. Proof of Theorem 3.1 (Trace-form representation)
Sketch of proof of Theorem 3.1.
We outline the argument and refer to Appendix A for the measure-theoretic details.
Fix a self-adjoint operator $D$ on $H$ and let $E_{D}(\cdot)$ be its spectral measure. Define the spectral trace measure $\nu_{D}$ on Borel sets $B\subset\mathbb{R}$ by
$\nu_{D}(B):=\operatorname{Tr}(E_{D}(B)).$
For bounded $B$ this is finite whenever $E_{D}(B)$ has finite rank; in general it may take the value $\infty$, but $E_{D}(B)$ is always a well-defined projection in the functional calculus of $D$.
Restrict $\mathcal{E}$ to the spectral projections of $D$ and set
$\mu_{D}(B):=\mathcal{E}(E_{D}(B)).$
By positivity, extensivity on orthogonal sums, projector-locality, and dominated continuity (applied to increasing sequences of projections dominated by a fixed spectral projection), the set function $B\mapsto\mu_{D}(B)$ extends from finite spectral partitions to a countably additive, finite measure on the $\sigma$-algebra generated by those spectral subsets, and $\mu_{D}(B)=0$ whenever $\nu_{D}(B)=0$. In particular, $\mu_{D}\ll\nu_{D}$; see Lemma A.7 and Remark A.8.
By the Radon-Nikodym theorem and the absolute continuity just noted, there exists a Borel-measurable function $w_{D}:\mathbb{R}\to[0,\infty)$ such that
$\mu_{D}(B)=\int_{B}w_{D}(\lambda)\,d\nu_{D}(\lambda)$
for all Borel sets $B\subset\mathbb{R}$ with $\nu_{D}(B)<\infty$. Thus, on the $\sigma$-finite part of the spectral trace measure
$\mathcal{A}_{D}:=\{B\subset\mathbb{R}\text{ Borel}:\nu_{D}(B)<\infty\},$
the evaluator on spectral projections admits a density $w_{D}$ with respect to $\nu_{D}$. We do not assume, and do not use, any Radon-Nikodym representation on Borel sets outside $\mathcal{A}_{D}$; all integrals involving $w_{D}$ are taken over sets of finite $\nu_{D}$-measure.
On the trace-class ideal $\mathcal{S}_{1}(H)$, the standard representation theorem for normal functionals on $B(H)$ (see Simon [3]*Chapter 1) yields a unique positive trace-class operator $T_{D}$ such that
$\mathcal{E}(A)=\operatorname{Tr}(T_{D}A)$
for all trace-class operators $A$ in the spectral algebra of $D$. In spectral terms this is equivalent to
$\tilde{\nu}_{D}(B):=\mu_{D}(B)=\operatorname{Tr}(T_{D}E_{D}(B))$
for all Borel sets $B\subset\mathbb{R}$ with $\nu_{D}(B)<\infty$. In particular, if $f(D)\in\mathcal{C}_{1}$, then
$\mathcal{E}(f(D))=\operatorname{Tr}\big(T_{D}f(D)\big),$
so on the trace-class envelope $\mathcal{C}_{1}$ the evaluator is a weighted trace with weight $T_{D}$ depending a priori on $D$. The Radon-Nikodym density $w_{D}$ and the trace-class operator $T_{D}$ provide two compatible scalar and operator-valued descriptions of the evaluator on the $\sigma$-finite part of the spectrum; neither is required for the construction of the global transform $h$.
Appendix A refines this per-operator description to a global transform that is independent of $D$. The finite-dimensional warm-up (Appendix A.1) shows that any evaluator satisfying the axioms must, on finite-dimensional self-adjoint operators, be given by a scalar function of eigenvalues:
$\mathcal{E}(X)=\operatorname{Tr}(h(X))$
for some function $h:\mathbb{R}\to\mathbb{R}$ determined by the values of $\mathcal{E}$ on rank-one spectral blocks. Using projector-locality and dominated continuity, Lemma A.2, Lemma A.3, and Lemma A.4 extend this construction to positive trace-class operators $X\in\mathcal{C}_{1}^{+}$ by approximating $X$ from below with finite-rank truncations and passing to the limit under domination.
The outcome (Lemma A.4) is a Borel-measurable, nondecreasing function $h:\mathbb{R}\to[0,\infty)$ such that
$\mathcal{E}(X)=\operatorname{Tr}\big(h(X)\big)$
for every $X\in\mathcal{C}_{1}^{+}$. Normalization and uniqueness (see Lemma A.5) then show that there exists a constant $c_{\mathcal{E}}>0$, independent of $D$ and $f$, such that
$\mathcal{E}(X)=c_{\mathcal{E}}\,\operatorname{Tr}\big(h(X)\big)$
for all positive $X\in\mathcal{C}_{1}$. Writing $X=f(D)$ with $f(D)\in\mathcal{C}_{1}^{+}$ yields the statement of Theorem 3.1 on the trace-class envelope:
$\mathcal{E}\big(f(D)\big)=c_{\mathcal{E}}\,\operatorname{Tr}\big(h(f(D))\big),$
with $h$ Borel-measurable and nondecreasing. The construction of $h$ and the proof of Theorem 3.1 on $\mathcal{C}_{1}$ are carried out entirely in Appendix A by the rank-one calibration and trace-class monotone convergence argument; the Radon-Nikodym density $w_{D}$ and the operator $T_{D}$ provide a measure-theoretic and trace-ideal interpretation of the evaluator on the $\sigma$-finite part of the spectrum but are not used as inputs to the existence or uniqueness of $h$. This completes the proof of Theorem 3.1, modulo the constructions and normalization arguments in Appendix A. ∎
4.2. Proof of Theorem 3.5 (Closure of spectral growth classes)
We keep the notation of Section 3.3. Throughout, $D$ and $E$ are self-adjoint operators with discrete spectra and counting functions $N_{D}$ and $N_{E}$ satisfying Convention 2.2. The counting-function reparametrization and tensor-product estimates we need are collected in Appendix B. In Section 4.2.1 we restate the monotone reparametrization lemma Lemma B.5 as Lemma 4.1. In Section 4.2.2 we prove the orthogonal-sum and tensor-product bounds; the polynomial tensor-product estimate is an immediate consequence of Lemma B.6 and Corollary B.8.
4.2.1. Monotone functional calculus
We first record the counting-function identity underlying part (1) of Theorem 3.5.
Lemma 4.1 (Monotone calculus bound).
Let $D$ be self-adjoint with discrete spectrum and counting function $N_{D}$, and let $f:[0,\infty)\to[0,\infty)$ be nondecreasing, unbounded, and regularly varying of positive index. Then
$N_{f(D)}(\Lambda)=N_{D}\bigl(f^{-1}(\Lambda)\bigr)$
for all $\Lambda\geq 0$, and if $N_{D}(\Lambda)\to\infty$ as $\Lambda\to\infty$ then
$N_{f(D)}(\Lambda)\sim N_{D}\bigl(f^{-1}(\Lambda)\bigr)\qquad(\Lambda\to\infty).$
Proof.
This is Lemma B.5 specialized to the setting of Section 3.3, with $f^{-1}$ as in Definition B.4. We recall the statement here for convenience and refer to Appendix B.2 for the proof. ∎
4.2.2. Orthogonal sum and tensor product
We now establish the behaviour of counting functions under orthogonal sum and tensor product. The orthogonal-sum part is elementary; the tensor-product estimate is a direct application of the convolution bounds developed in Appendix B.3.
Lemma 4.2 (Orthogonal sum and tensor product).
Let $D$ and $E$ be self-adjoint with discrete spectra and counting functions $N_{D},N_{E}$.
1. For all $\Lambda\geq 0$,
$N_{D\oplus E}(\Lambda)=N_{D}(\Lambda)+N_{E}(\Lambda).$
2. Suppose there exist exponents $d_{D},d_{E}>0$ and, for every $\varepsilon>0$, constants $C_{D,\varepsilon},C_{E,\varepsilon}>0$ such that
$N_{D}(\Lambda)\leq C_{D,\varepsilon}\,\Lambda^{d_{D}+\varepsilon},\qquad N_{E}(\Lambda)\leq C_{E,\varepsilon}\,\Lambda^{d_{E}+\varepsilon}$
for all sufficiently large $\Lambda$. Then for every $\varepsilon>0$ there exists $C_{\varepsilon}>0$ such that
$N_{D\otimes E}(\Lambda)\leq C_{\varepsilon}\,\Lambda^{d_{D}+d_{E}+\varepsilon}$
for all sufficiently large $\Lambda$. In particular, if $D,E\in\mathcal{G}_{\mathrm{poly}}$ then $D\otimes E\in\mathcal{G}_{\mathrm{poly}}$.
Proof.
(1) The spectrum of $D\oplus E$ is the multiset union of the spectra of $D$ and $E$, so eigenvalues and multiplicities simply add. Hence
$N_{D\oplus E}(\Lambda)=\#\{n:|\lambda_{n}(D)|\leq\Lambda\}+\#\{m:|\lambda_{m}(E)|\leq\Lambda\}=N_{D}(\Lambda)+N_{E}(\Lambda).$
(2) Write $\{a_{j}\}_{j\geq 1}$ and $\{b_{k}\}_{k\geq 1}$ for the eigenvalues of $D$ and $E$, counted with multiplicity and ordered so that $|a_{j}|$ and $|b_{k}|$ are nondecreasing. The spectrum of $D\otimes E$ consists of all products $a_{j}b_{k}$, counted with multiplicity, and
$N_{D\otimes E}(\Lambda)=\#\{(j,k):|a_{j}b_{k}|\leq\Lambda\}=\sum_{j\geq 1}N_{E}\bigl(\Lambda/|a_{j}|\bigr).$
Fix $\varepsilon>0$ and set $\delta=\varepsilon/2$. By hypothesis there exist constants $C_{D},C_{E}>0$ such that
$N_{D}(t)\leq C_{D}\,t^{d_{D}+\delta},\qquad N_{E}(t)\leq C_{E}\,t^{d_{E}+\delta}$
for all sufficiently large $t$. These growth bounds are exactly the hypotheses of Appendix B.3. In particular, Lemma B.6 and Corollary B.8 applied with these exponents give
$N_{D\otimes E}(\Lambda)\leq C^{\prime\prime}_{\delta}\,\Lambda^{d_{D}+d_{E}+2\delta}$
for all sufficiently large $\Lambda$. Since $2\delta=\varepsilon$, we may absorb $C^{\prime\prime}_{\delta}$ into a constant $C_{\varepsilon}$ and obtain
$N_{D\otimes E}(\Lambda)\leq C_{\varepsilon}\,\Lambda^{d_{D}+d_{E}+\varepsilon}$
for all sufficiently large $\Lambda$, which is precisely the polynomial tensor-product estimate stated in part (2). ∎
4.2.3. Conclusion of the proof
We now prove Theorem 3.5.
Proof of Theorem 3.5.
The disjointness of the growth classes $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\mathrm{log}}$ follows from the incompatibility of their asymptotic forms. For example, if $N_{D}(\Lambda)\sim C\,\Lambda^{d}$ with $C,d>0$ and also $N_{D}(\Lambda)\sim\exp(c\,\Lambda^{\alpha})$ with $c>0$ and $0<\alpha<1$, then comparing logarithms as $\Lambda\to\infty$ yields a contradiction. More generally, the classification of regularly varying functions shows that polynomial, stretched-exponential, and logarithmic-or-slower growth regimes are mutually exclusive; see Appendix B and Bingham-Goldie-Teugels [5]*Theorem 1.3.1 and Section 1.4.
For part (1), let $f$ be as in Theorem 3.5(1) and assume $D\in\mathcal{G}_{\mathrm{poly}}$ with
$N_{D}(\Lambda)\sim C\,\Lambda^{d}\qquad(\Lambda\to\infty).$
By Lemma 4.1 we have
$N_{f(D)}(\Lambda)=N_{D}\bigl(f^{-1}(\Lambda)\bigr).$
Since $f$ is regularly varying of positive index, so is its generalized inverse $f^{-1}$ on large arguments, and the asymptotic behaviour of $N_{D}$ transfers along composition with $f^{-1}$. In particular, $N_{f(D)}(\Lambda)$ again has polynomial growth, so $f(D)\in\mathcal{G}_{\mathrm{poly}}$.
For part (2), Lemma 4.2(1) shows that if $D,E$ lie in the same growth class, then
$N_{D\oplus E}(\Lambda)=N_{D}(\Lambda)+N_{E}(\Lambda)$
inherits the same asymptotic form, so the corresponding family is closed under finite orthogonal sums.
For part (3), Lemma 4.2(2) gives the polynomial tensor-product bound. If $D,E\in\mathcal{G}_{\mathrm{poly}}$, then the hypotheses of Lemma 4.2(2) are satisfied for some exponents $d_{D},d_{E}$, and we conclude that
$N_{D\otimes E}(\Lambda)=O\bigl(\Lambda^{d_{D}+d_{E}+\varepsilon}\bigr)$
for every $\varepsilon>0$, so $D\otimes E\in\mathcal{G}_{\mathrm{poly}}$.
Combining these three parts, we obtain the closure properties asserted in Theorem 3.5. ∎
5. Examples
This section collects representative examples illustrating the evaluator framework and the trace-form representation theorem in concrete spectral geometries. In each case, a self-adjoint operator $D$ provides a model spectral geometry through its bounded Borel transforms $f(D)$, and the associated values $\mathcal{E}(f(D))$ are computed via the trace-form representation on $\mathcal{C}_{1}$. The examples are chosen to cover the basic growth classes (polynomial, stretched-exponential, and logarithmic-or-slower), as well as to emphasize how dominated continuity and cutoff procedures act on spectral geometries rather than on scalar data.
Taken together, these examples show how the general calculus specializes in concrete settings. Once the spectral geometry is fixed, every evaluator satisfying Axioms A1-A5 assigns values by integrating a single profile $h$ against the spectral measure, and the growth class of $D$ determines which transforms $f(D)$ fall into the trace-class envelope $\mathcal{C}_{1}$. In this sense the examples are not separate constructions but instances of the same mechanism: lawful scalar data on these models are fully dictated by spectral geometry and the fixed profile $h$, as encoded in the trace-form representation of Theorem 3.1.
5.1. Finite-rank diagonal operators
Let $H=\ell^{2}(\mathbb{N})$ and consider a finite-rank diagonal operator
$D=\operatorname{diag}(\lambda_{1},\dots,\lambda_{m},0,0,\dots),$
with real eigenvalues $\lambda_{1},\dots,\lambda_{m}$. For any bounded continuous function $f:\mathbb{R}\to\mathbb{R}$, we have
$f(D)=\operatorname{diag}\big(f(\lambda_{1}),\dots,f(\lambda_{m}),f(0),f(0),\dots\big),$
and the eigenvalues of $f(D)$ are simply $\{f(\lambda_{j})\}_{j\geq 1}$ with multiplicity.
If $f(D)$ is trace class, then
$\operatorname{Tr}(f(D))=\sum_{j=1}^{\infty}f(\lambda_{j})=\sum_{j=1}^{m}f(\lambda_{j})+\sum_{j>m}f(0),$
and the second sum vanishes whenever $f(0)=0$. In particular, for finite-rank $D$ and bounded continuous $f$ with $f(0)=0$, the trace is a finite sum of eigenvalues.
Under the hypotheses of Theorem 3.1, the evaluator $\mathcal{E}$ restricted to this finite-rank class must agree with a weighted spectral trace:
$\mathcal{E}(f(D))=c_{\mathcal{E}}\sum_{j=1}^{m}h(f(\lambda_{j})),$
for some Borel-measurable nondecreasing $h$. This exhibits the trace-form representation in the simplest setting and shows that, on finite-rank diagonals, $\mathcal{E}$ is determined entirely by the eigenvalues of the input.
5.2. Diagonal models for growth classes
We next give basic diagonal examples realizing each of the growth classes introduced in Definition 3.4.
5.2.1. Polynomial growth
Set
$D_{\mathrm{poly}}=\operatorname{diag}(1,2,3,\dots)$
on $\ell^{2}(\mathbb{N})$. Then the counting function
$N_{D_{\mathrm{poly}}}(\Lambda)=\#\{n\geq 1:n\leq\Lambda\}=\lfloor\Lambda\rfloor$
satisfies
$N_{D_{\mathrm{poly}}}(\Lambda)\sim\Lambda\quad(\Lambda\to\infty),$
so $D_{\mathrm{poly}}\in\mathcal{G}_{\mathrm{poly}}$ with exponent $d=1$. Note that $D_{\mathrm{poly}}$ is not trace class, hence $D_{\mathrm{poly}}\notin\mathcal{C}_{1}$; it is used here purely as a growth model and not as an input for the trace-form representation of Theorem 3.1.
As a simple test of the monotone functional calculus in the polynomial regime, fix $\beta>0$ and set $f(t)=t^{\beta}$ on $0,\infty)$. Then
$f(D_{\mathrm{poly}})=\operatorname{diag}(n^{\beta})_{n\geq 1},$
and
$N_{f(D_{\mathrm{poly}})}(\Lambda)=\#\{n\geq 1:n^{\beta}\leq\Lambda\}=\#\{n\geq 1:n\leq\Lambda^{1/\beta}\}\sim\Lambda^{1/\beta}.$
In particular $f(D_{\mathrm{poly}})\in\mathcal{G}_{\mathrm{poly}}$, in agreement with Theorem 3.5(1).
5.2.2. Stretched-exponential behavior
Fix constants $c>0$ and $0<\alpha<1$, and define
$D_{\mathrm{str}}=\operatorname{diag}\Big(\Big(\tfrac{1}{c}\log(n+1)\Big)^{1/\alpha}\Big)_{n\geq 1}.$
Then for $\Lambda\geq 0$,
$N_{D_{\mathrm{str}}}(\Lambda)=\#\Big\{n\geq 1:\Big(\tfrac{1}{c}\log(n+1)\Big)^{1/\alpha}\leq\Lambda\Big\}=\#\{n\geq 1:n+1\leq\exp(c\,\Lambda^{\alpha})\}.$
Hence
$N_{D_{\mathrm{str}}}(\Lambda)\sim\exp(c\,\Lambda^{\alpha})\quad(\Lambda\to\infty),$
so $D_{\mathrm{str}}\in\mathcal{G}_{\mathrm{str}}$. Note that “stretched-exponential” refers to the asymptotic growth of the counting function $N_{D}$, not to the eigenvalue sequence $(\lambda_{n}(D))$ itself.
5.2.3. Logarithmic or slower growth
Consider
$D_{\log}=\operatorname{diag}\big(e^{n}\big)_{n\geq 1}.$
Then
$N_{D_{\log}}(\Lambda)=\#\{n\geq 1:e^{n}\leq\Lambda\}=\#\{n\geq 1:n\leq\log\Lambda\}\sim\log\Lambda$
as $\Lambda\to\infty$. Thus $D_{\log}$ has logarithmic spectral growth in the sense of Definition 3.4(3) and provides a basic model for the “logarithmic or slower” class $\mathcal{G}_{\mathrm{log}}$.
5.3. Evaluators on Laplacian-type spectra
Let $G$ be a finite connected graph with vertex set $V$ and combinatorial Laplacian $L$ acting on $\ell^{2}(V)$. The operator $L$ is self-adjoint and positive, with purely discrete spectrum
$0=\lambda_{1}\leq\lambda_{2}\leq\dots\leq\lambda_{|V|}.$
For any bounded continuous $f:\mathbb{R}_{+}\to\mathbb{R}$ with $f(0)=0$, the transform $f(L)$ is trace class and
$\operatorname{Tr}(f(L))=\sum_{j=2}^{|V|}f(\lambda_{j})$ (5.1)
Under the standing assumptions, Theorem 3.1 implies that any evaluator $\mathcal{E}$ on this class agrees with a weighted trace-form observable:
$\mathcal{E}(f(L))=c_{\mathcal{E}}\sum_{j=2}^{|V|}h(f(\lambda_{j}))$ (5.2)
for some Borel-measurable nondecreasing $h$. In particular, choosing $h(t)=t$ and $c_{\mathcal{E}}=1$ recovers the usual trace as a special case.
On compact Riemannian manifolds, Laplace-type operators exhibit polynomial growth of eigenvalues governed by Weyl-law asymptotics, and the trace-form representation continues to apply under the same structural hypotheses. These examples show that the abstract evaluator calculus is compatible with standard spectral-geometry models.
5.4. Stability under sums and tensor products
We illustrate the closure statements in Theorem 3.5 with simple constructions.
5.4.1. Orthogonal sums
Let $D_{\mathrm{poly}}$ be as in Section 5.2, and let $E$ be another diagonal operator with the same growth rate, for instance
$E=\operatorname{diag}(2,4,6,\dots).$
Then for each $\Lambda$,
$N_{D_{\mathrm{poly}}\oplus E}(\Lambda)=N_{D_{\mathrm{poly}}}(\Lambda)+N_{E}(\Lambda)\sim 2\Lambda,$
so the orthogonal sum remains in $\mathcal{G}_{\mathrm{poly}}$ with the same exponent.
5.4.2. Tensor products
Let $D_{\mathrm{poly}}$ be as above and consider the tensor product $D_{\mathrm{poly}}\otimes D_{\mathrm{poly}}$ acting on $\ell^{2}(\mathbb{N}\times\mathbb{N})$. The eigenvalues of $D_{\mathrm{poly}}\otimes D_{\mathrm{poly}}$ are $\{mn:m,n\geq 1\}$, counted with multiplicity. The associated counting function satisfies
$N_{D_{\mathrm{poly}}\otimes D_{\mathrm{poly}}}(\Lambda)=\#\{(m,n)\in\mathbb{N}^{2}:mn\leq\Lambda\}\leq C\,\Lambda^{1+\varepsilon}$
for any $\varepsilon>0$ and some $C>0$, so $D_{\mathrm{poly}}\otimes D_{\mathrm{poly}}$ lies in $\mathcal{G}_{\mathrm{poly}}$ with exponent at most $2$. This is consistent with the tensor-product estimate in Lemma 4.2.
5.4.3. Monotone transforms
If we apply a monotone transform $f(\lambda)=\lambda^{\gamma}$ with $\gamma>0$ to $D_{\mathrm{poly}}$, we obtain
$f(D_{\mathrm{poly}})=\operatorname{diag}(1^{\gamma},2^{\gamma},3^{\gamma},\dots),$
and the counting function satisfies
$N_{f(D_{\mathrm{poly}})}(\Lambda)=\#\{n\geq 1:n^{\gamma}\leq\Lambda\}\sim\Lambda^{1/\gamma},$
so polynomial growth is preserved under such transforms. This is a concrete instance of the monotone calculus bound in Lemma 4.1.
These examples confirm, in explicit models, that the growth classes and the structural operations of the spectral calculus are compatible in the way asserted by Theorem 3.5.
5.5. Evaluator limits under dominated convergence
Finally, we illustrate the dominated-continuity aspect of the evaluator calculus.
Let $D$ be a fixed self-adjoint operator with discrete spectrum and counting function $N_{D}$. Let $f:\mathbb{R}\to\mathbb{R}$ be bounded and continuous, and for each $k\geq 1$ define
$f_{k}(\lambda):=f(\lambda)\,\mathbf{1}_{[-k,k]}(\lambda),\qquad\lambda\in\mathbb{R}.$
Then $f_{k}\to f$ uniformly and
$|f_{k}(\lambda)|\leq\|f\|_{\infty}\quad\text{for all }\lambda\in\mathbb{R},\ k\geq 1.$
By the spectral theorem, $f_{k}(D)\to f(D)$ in operator norm, hence strongly. By stability of $\mathcal{C}$, each $f_{k}(D)$ and $f(D)$ lies in $\mathcal{C}$.
Note that
$|f_{k}(D)|\leq\|f\|_{\infty}I\quad\text{for all }k\geq 1.$
Set
$Y:=\|f\|_{\infty}I.$
By the normalization axiom, $I\in\mathcal{C}$ and $\mathcal{E}(I)\in(0,\infty)$, so $Y\in\mathcal{C}$ and
$\mathcal{E}(Y)=\|f\|_{\infty}\,\mathcal{E}(I)<\infty.$
Thus the sequence $\{f_{k}(D)\}$ is dominated by a spectral transform $Y$ with finite evaluation, and the dominated continuity axiom applies. We obtain
$\mathcal{E}(f_{k}(D))\to\mathcal{E}(f(D)).$
Assume in addition that each $f_{k}(D)$ and $f(D)$ lies in the trace-class envelope $\mathcal{C}_{1}$. Then Theorem 3.1 gives
$\mathcal{E}(f_{k}(D))=c_{\mathcal{E}}\,\operatorname{Tr}\big(h(f_{k}(D))\big),$
for the Borel-measurable nondecreasing transform $h$ and constant $c_{\mathcal{E}}>0$ provided by Theorem 3.1. Under the same domination hypotheses, dominated convergence on the spectral side yields
$\operatorname{Tr}\big(h(f_{k}(D))\big)\to\operatorname{Tr}\big(h(f(D))\big),$
so
$\mathcal{E}(f(D))=c_{\mathcal{E}}\,\operatorname{Tr}\big(h(f(D))\big).$
This example shows how dominated continuity at the operator level is mirrored by dominated convergence at the spectral level and how the trace-form representation on $\mathcal{C}_{1}$ remains stable under such limiting procedures. It also illustrates the type of hypotheses under which evaluator limits and trace-form limits commute. In particular, this construction realizes the dominated-continuity axiom A4 in a concrete setting: the truncations $f_{k}(D)$ form a monotone sequence converging strongly to $f(D)$, are dominated by the spectral transform $Y=\|f\|_{\infty}I$ with $\mathcal{E}(Y)<\infty$, and therefore provide an explicit example in which the hypotheses and conclusion of Axiom A4 are all satisfied.
6. Consequences, limitations, and falsifiability
6.1. Consequences
The main theorems have several direct consequences for the structure of lawful evaluators and for the behavior of spectral growth classes. On the one hand, the trace-form representation theorem reduces any evaluator satisfying Axioms A1-A5 to a weighted spectral trace on the trace-class envelope $\mathcal{C}_{1}$, with no additional scalar degrees of freedom. On the other hand, the growth taxonomy organizes self-adjoint operators into regimes in which trace-form evaluations are well defined and stable under the basic constructions of the calculus. We record a few formal consequences in this direction.
Corollary 6.1 (Trace-form reduction on the trace-class envelope).
Let $\mathcal{C}$ be a stable class of spectral transforms and $\mathcal{E}:\mathcal{C}\to[0,\infty]$ an evaluator satisfying the standing assumptions of Section 3.1. Write $\mathcal{C}_{1}=\mathcal{C}\cap\mathcal{S}_{1}(H)$ for the trace-class envelope. Then for every bounded continuous $f$ and self-adjoint $D$ such that $f(D)\in\mathcal{C}_{1}$,
$\mathcal{E}(f(D))=c_{\mathcal{E}}\,\operatorname{Tr}\big(h(f(D))\big),$
for some Borel-measurable nondecreasing $h:\mathbb{R}\to[0,\infty)$ and $c_{\mathcal{E}}>0$ independent of $D$ and $f$. In particular, on the subclass of trace-class operators in $\mathcal{C}$, the evaluator is a weighted spectral trace.
Proof.
This is an immediate consequence of Theorem 3.1 and the construction described in Section 4.1. ∎
Corollary 6.2 (Spectral growth families: closure and taxonomy).
Let $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\mathrm{log}}$ denote the polynomial, stretched-exponential, and logarithmic-or-slower growth classes defined in Definition 3.4. Then:
1. The families $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\mathrm{log}}$ are pairwise disjoint. 2. Each family is closed under monotone functional calculus by regularly varying transforms. 3. Each family is closed under finite orthogonal sums. 4. In the polynomial case, each family is closed under tensor products: if $D,E\in\mathcal{G}_{\mathrm{poly}}$, then $D\otimes E\in\mathcal{G}_{\mathrm{poly}}$.
These closure properties ensure that the growth taxonomy is stable under the basic structural operations of the spectral calculus.
Proof.
This follows directly from Theorem 3.5 and the counting-function estimates established in Section 4.2. ∎
Corollary 6.3 (Evaluator continuity under dominated convergence).
Let $D$ be self-adjoint and $\{f_{k}\}$ a sequence of bounded continuous functions on $\mathbb{R}$ converging uniformly to a bounded continuous function $f$. Suppose $f_{k}(D),f(D)\in\mathcal{C}$ for all $k$ and there exists a positive spectral transform $Y\in\mathcal{C}$ with $\mathcal{E}(Y)<\infty$ such that
$|f_{k}(D)|\leq Y\quad\text{for all }k.$
Then
$\mathcal{E}(f_{k}(D))\to\mathcal{E}(f(D)),$
and, whenever the trace-form representation applies (i.e. when $f_{k}(D),f(D)\in\mathcal{C}_{1}$),
$\operatorname{Tr}\big(h(f_{k}(D))\big)\to\operatorname{Tr}\big(h(f(D))\big).$
In particular, evaluator limits and trace-form limits commute under the same domination hypotheses. This provides a basic legality criterion for passing to limits in families of spectral transforms.
Proof.
This follows from the dominated continuity assumption of Section 3.1 applied to the trace-form representation in Theorem 3.1. ∎
6.2. Limitations
The hypotheses used in the main theorems are close to minimal, but they impose several intrinsic limitations on the framework. It is useful to separate two global structural constraints from a set of more technical restrictions.
  • Spectral representation (SOC-I).
    SOC-I is formulated in the setting of the spectral theorem: each model is a self-adjoint operator D on a separable Hilbert space together with its projection-valued spectral measure, and evaluators act on bounded Borel transforms f(D). Other operator classes require different functional-calculus inputs and, crucially, a different realization of the localization (coarse-graining) gate; the sectorial/holomorphic model case is developed separately.
  • Spectral rigidity on $\mathcal{C}_{1}$.
    Theorem 3.1 (trace-form representation) is proved only for inputs in the trace-class envelope $\mathcal{C}_{1}:=\mathcal{C}\cap\mathcal{S}_{1}(H)$. Within $\mathcal{C}_{1}$, the axioms A1-A5 force any evaluator to be of trace form with a single Borel nondecreasing profile $h$ and an overall normalization $c_{\mathcal{E}}$; in particular, there is no additional scalar freedom once these choices are fixed. Outside $\mathcal{C}_{1}$ the calculus does not assert trace-form representation, and additional hypotheses would be needed to control evaluators on non-trace-class transforms.
In addition, the following technical limitations are built into the present formulation.
  1. 1.
    Dependence on projector-locality (PVM-localized model case).
    Theorem 3.1 (trace-form representation) relies critically on projector-locality, i.e. additivity across spectral (PVM) partitions. If $\mathcal{E}$ is only unitary invariant and extensive on orthogonal sums, but not projector-local, then there exist natural examples (e.g. operator norm on finite-rank self-adjoint operators) that cannot be written in trace form and violate the additivity required on spectral partitions. Such counterexamples are discussed in detail in Appendix C. In the sectorial model case, this axiom is replaced by a scale-band locality principle rather than by PVM additivity.
  2. 2.
    Dependence on dominated continuity.
    Dropping the dominated continuity requirement allows construction of evaluators that are finitely additive on spectral projections but fail to be countably additive, breaking the Radon-Nikodym representation used in Section 4.1. In particular, it is possible to arrange sequences of spectral cutoffs $f_{k}(D)$ for which $f_{k}\to f$ pointwise and $\{f_{k}(D)\}$ is uniformly bounded, yet $\mathcal{E}(f_{k}(D))$ does not converge to $\mathcal{E}(f(D))$. Explicit counterexamples appear in Appendix C.
  3. 3.
    Discrete-spectrum hypothesis for growth.
    The spectral growth taxonomy in Definition 3.4 and Theorem 3.5 assumes that $D$ has discrete spectrum with finite multiplicities and that the counting function $N_{D}(\Lambda)$ satisfies the regularity conditions of Convention 2.2. For operators with continuous spectrum or highly degenerate spectral measures, a different notion of growth (e.g. via spectral density functions or integrated density of states) is required. These extensions are not treated here.
  4. 4.
    Regular variation assumptions.
    The closure of growth classes under monotone functional calculus uses regular variation for the transforming function $f$. If $f$ has oscillatory or highly irregular behavior at infinity, the simple asymptotic relations of Lemma 4.1 may fail, and the effect on growth class can be subtler. In such cases, additional hypotheses on $f$ are needed.
  5. 5.
    Scope of regularization.
    Regularized traces are considered only when they arise from explicit spectral truncation or weighting compatible with the evaluator axioms. More aggressive renormalization schemes, or those relying on analytic continuation in external parameters, are beyond the scope of the present framework.
These limitations mark the boundary of the calculus developed in this paper and indicate where additional hypotheses or different techniques would be required.
6.3. Falsifiable internal tests
The framework presented here makes several concrete claims about evaluators on spectral geometries and about the interaction between value assignments and spectral growth. In particular, it asserts that within the trace-class envelope $\mathcal{C}_{1}$ every lawful evaluator is of trace form with a nondecreasing profile, and that the basic growth classes are stable under the standard constructions of the spectral calculus. These claims are intrinsically falsifiable: they can be tested by constructing explicit operators and evaluators and checking whether the predicted trace-form behavior and growth stability hold. We record a few representative internal tests.
Test F1 (Trace-form representation on $\mathcal{C}_{1}$).
Setup: Fix a stable class $\mathcal{C}$ of spectral transforms and an evaluator $\mathcal{E}$ satisfying the axioms of Section 3.1, and write $\mathcal{C}_{1}=\mathcal{C}\cap\mathcal{S}_{1}(H)$.
Claim: There exist a Borel-measurable nondecreasing function $h:\mathbb{R}\to[0,\infty)$ and a constant $c_{\mathcal{E}}>0$ such that for every bounded continuous $f$ and self-adjoint $D$ with $f(D)\in\mathcal{C}_{1}$,
$\mathcal{E}(f(D))=c_{\mathcal{E}}\,\operatorname{Tr}\big(h(f(D))\big).$
In particular, once $h$ and $c_{\mathcal{E}}$ are fixed, all values of $\mathcal{E}$ on $\mathcal{C}_{1}$ are determined by the spectra of the underlying operators.
Refutation strategy: Exhibit a specific $\mathcal{C}$, an evaluator $\mathcal{E}$, and a family of test pairs $(f,D)$ with $f(D)\in\mathcal{C}_{1}$ for which no choice of Borel nondecreasing $h$ and scalar $c_{\mathcal{E}}$ can satisfy the trace-form identity above. Any such example would contradict Theorem 3.1 and falsify the trace-form representation claim for that evaluator.
Test F2 (Growth family closure).
Setup: Let $D$ lie in one of the growth classes $\mathcal{G}_{\mathrm{poly}},\mathcal{G}_{\mathrm{str}},\mathcal{G}_{\mathrm{log}}$. Consider monotone regularly varying transforms $f$, orthogonal sums $D\oplus E$, and tensor products $D\otimes E$ with partners $E$ in the same class.
Claim: The transformed operators remain in the same growth class and satisfy the asymptotic bounds stated in Theorem 3.5.
Refutation strategy: Produce explicit $D,E,f$ with well-understood spectra for which the resulting $N_{f(D)}$, $N_{D\oplus E}$, or $N_{D\otimes E}$ leaves the claimed growth class.
Test F3 (Evaluator continuity under dominated limits).
Setup: Fix a self-adjoint $D$, a sequence of bounded continuous functions $f_{k}\to f$ uniformly, and a dominating spectral transform $Y$ as in Corollary 6.3.
Claim: The evaluator satisfies
$\mathcal{E}(f_{k}(D))\to\mathcal{E}(f(D)).$
Refutation strategy: Produce an evaluator $\mathcal{E}$ and data $D,f_{k},f,Y$ satisfying all structural assumptions for which the limit fails to exist or to equal $\mathcal{E}(f(D))$.
Test F4 (Stability of discrete models under interpolation).
Setup: Let $\{x_{n}\}_{n\geq 1}$ and $\{y_{n}\}_{n\geq 1}$ be two real sequences of bounded variation, and let $D_{x},D_{y}$ be the associated diagonal operators. Consider families of interpolants $F_{x,N},F_{y,N}$ that match $x_{n},y_{n}$ on integer points and are of uniformly bounded variation on compact intervals.
Claim: If the induced evaluators $\mathcal{E}(f(D_{x}))$ and $\mathcal{E}(f(D_{y}))$ agree for all bounded continuous $f$ in a suitable class, then the limiting interpolants $F_{x},F_{y}$ coincide almost everywhere with respect to the spectral trace measures of $D_{x},D_{y}$.
Refutation strategy: Construct discrete models $\{x_{n}\},\{y_{n}\}$ and an evaluator $\mathcal{E}$ satisfying the axioms such that equality of all evaluated transforms fails to force equality of the corresponding interpolants.
These tests illustrate how the abstract statements of the calculus can, in principle, be challenged within explicit operator models. Any successful refutation along these lines would pinpoint a gap in the hypotheses or a genuine limitation of the framework.
7. Conclusion
We have developed a spectral-operator calculus on separable Hilbert space based on a small collection of structural axioms for evaluators acting on spectral transforms. Under unitary invariance, projector-locality, dominated continuity, and stability of the underlying class, every extensive evaluator is forced to be of trace form: it coincides with the trace of a Borel-measurable nondecreasing transform of the input operator, up to normalization. This establishes a rigidity principle for spectral evaluation: once the structural requirements are fixed, no essentially new extensive evaluators exist beyond trace-type functionals.
Alongside evaluator rigidity, we introduced a spectral growth taxonomy for self-adjoint operators with discrete spectrum, organized by counting-function asymptotics. Polynomial, stretched-exponential, and logarithmic-or-slower growth classes were shown to be disjoint and stable under the basic operations of the spectral calculus: monotone functional calculus by regularly varying transforms, finite orthogonal sums, and (in the polynomial case) tensor products. These closure properties ensure that growth classes behave predictably under the structural manipulations needed for kernel constructions and flow limits.
A recurring theme has been the role of dominated convergence at both the operator and spectral levels. The dominated continuity axiom for the evaluator enables a Radon-Nikodym representation on spectral projections and justifies passing to limits under spectral truncations and reparametrizations. In model classes, this manifests as the commutation of evaluator limits with trace-form limits under natural domination hypotheses, providing a legal framework for working with families of spectral transforms and cutoff procedures.
The limitations recorded in Section 6.2 indicate where further hypotheses or different techniques would be required: relaxing projector-locality, dropping dominated continuity, treating continuous or highly degenerate spectra, or employing more aggressive regularizations lie beyond the scope of the present framework. Within its stated domain, however, the calculus developed here supplies a robust backbone for subsequent work. Later papers in this series take the trace-form evaluator and growth taxonomy as given and build on them to analyze admissible kernels and cutoff schemes, analytic continuation procedures, and boundary and reconstruction problems in more elaborate operator-theoretic settings, including sectorial/holomorphic models where localization is implemented on log-scale rather than by spectral projections.
Acknowledgements
The author acknowledges the use of GPT models from OpenAI and Claude models from Anthropic for assistance in outlining, literature survey, refining formal statements, and generating illustrative examples. All mathematical results, interpretations, and final decisions regarding this manuscript are the responsibility of the author.
Appendix A Proof of Theorem 3.1 (Trace-form representation)
Throughout this appendix we work under the standing assumptions of Section 3.1. In particular:
  • $\mathcal{C}$ is a stable class of spectral transforms, closed under bounded Borel functional calculus, finite orthogonal sums, and monotone strong limits;
  • $\mathcal{E}:\mathcal{C}\to[0,\infty]$ is unitary invariant, extensive on orthogonal sums, projector-local, and satisfies dominated continuity with $\mathcal{E}(I)\in(0,\infty)$.
We write
$\mathcal{C}_{1}:=\mathcal{C}\cap\mathcal{S}_{1}(H)$
for the trace-class envelope of $\mathcal{C}$, and
$\mathcal{C}_{1}^{+}:=\{X\in\mathcal{C}_{1}:X\geq 0\}$
for its positive cone. We restrict attention to inputs $X\in\mathcal{C}_{1}^{+}$. Extension to general self-adjoint trace-class operators follows by decomposing into positive and negative parts.
Our goal is to prove Theorem 3.1: on $\mathcal{C}_{1}^{+}$ the evaluator $\mathcal{E}$ is of trace form with a Borel-measurable nondecreasing transform $h$, unique up to a positive scalar factor.
A.1. Finite-dimensional warm-up
We begin with the finite-dimensional case, which already contains the essential structure.
Let $H=\mathbb{C}^{n}$ and let $\mathcal{C}_{\mathrm{fin}}$ be the set of all self-adjoint $n\times n$ matrices. Let $\mathcal{E}_{\mathrm{fin}}:\mathcal{C}_{\mathrm{fin}}\to\mathbb{R}$ satisfy:
  1. 1.
    Unitary invariance. $\mathcal{E}_{\mathrm{fin}}(U^{*}XU)=\mathcal{E}_{\mathrm{fin}}(X)$ for all unitaries $U$.
  2. 2.
    Extensivity. If $X,Y$ act on orthogonal subspaces and $X\oplus Y$ is their block-diagonal sum, then
    $\mathcal{E}_{\mathrm{fin}}(X\oplus Y)=\mathcal{E}_{\mathrm{fin}}(X)+\mathcal{E}_{\mathrm{fin}}(Y).$
  3. 3.
    Projector-locality. If $X$ has a spectral decomposition
    $X=\sum_{j=1}^{m}\lambda_{j}P_{j}$
    with pairwise orthogonal projections $P_{j}$, then
    $\mathcal{E}_{\mathrm{fin}}(X)=\sum_{j=1}^{m}\mathcal{E}_{\mathrm{fin}}(\lambda_{j}P_{j}).$
  4. 4.
    Normalization. There exists a nonzero projection $P_{0}$ with $\mathcal{E}_{\mathrm{fin}}(P_{0})\in(0,\infty)$.
We claim that $\mathcal{E}_{\mathrm{fin}}$ is necessarily of trace form.
Lemma A.1 (Finite-dimensional representation).
There exist a function $h:\mathbb{R}\to\mathbb{R}$ and a constant $c_{\mathcal{E}}>0$ such that
$\mathcal{E}_{\mathrm{fin}}(X)=c_{\mathcal{E}}\;\operatorname{Tr}\big(h(X)\big)$
for all self-adjoint $X\in\mathcal{C}_{\mathrm{fin}}$, where $h(X)$ is defined by functional calculus on the eigenvalues of $X$.
Proof.
Fix a rank-one projection $P$. By unitary invariance, $\mathcal{E}_{\mathrm{fin}}(\lambda P)$ depends only on $\lambda$ and not on the particular choice of $P$. Define
$h(\lambda):=\mathcal{E}_{\mathrm{fin}}(\lambda P),$
so that $h:\mathbb{R}\to\mathbb{R}$ is well-defined.
If $Q$ is any finite-rank projection of rank $r$, we can write $Q=P_{1}\oplus\cdots\oplus P_{r}$ as a sum of rank-one projections. By extensivity and unitary invariance,
$\mathcal{E}_{\mathrm{fin}}(\lambda Q)=\sum_{j=1}^{r}\mathcal{E}_{\mathrm{fin}}(\lambda P_{j})=r\,h(\lambda)=\operatorname{Tr}\big(h(\lambda)Q\big).$
Now let $X$ be an arbitrary self-adjoint matrix with spectral decomposition
$X=\sum_{j=1}^{m}\lambda_{j}P_{j},$
where the $P_{j}$ are pairwise orthogonal projections. By projector-locality and the above calculation,
$\mathcal{E}_{\mathrm{fin}}(X)=\sum_{j=1}^{m}\mathcal{E}_{\mathrm{fin}}(\lambda_{j}P_{j})=\sum_{j=1}^{m}\operatorname{Tr}\big(h(\lambda_{j})P_{j}\big)=\operatorname{Tr}\Big(\sum_{j=1}^{m}h(\lambda_{j})P_{j}\Big)=\operatorname{Tr}\big(h(X)\big).$
Finally, choose a nonzero projection $P_{0}$ as in the normalization assumption and set
$c_{\mathcal{E}}:=\frac{\mathcal{E}_{\mathrm{fin}}(P_{0})}{\operatorname{Tr}(h(P_{0}))}.$
Then
$\mathcal{E}_{\mathrm{fin}}(X)=\operatorname{Tr}\big(h(X)\big)=c_{\mathcal{E}}\;\frac{1}{c_{\mathcal{E}}}\,\operatorname{Tr}\big(h(X)\big)=c_{\mathcal{E}}\;\operatorname{Tr}\big(\tilde{h}(X)\big),$
where $\tilde{h}=h/c_{\mathcal{E}}$ absorbs the normalization. Renaming $\tilde{h}$ as $h$ yields the desired form. Uniqueness of $h$ up to a positive scalar factor follows by evaluating on rank-one projections and comparing the resulting scalar functions. ∎
The infinite-dimensional proof mirrors this structure, but we must take care with limits and trace-class inputs. We now construct the global transform $h$ and the trace-form representation on $\mathcal{C}_{1}^{+}$.
A.2. Construction of the transform $h$ on trace-class inputs
We work on a separable Hilbert space $H$ and restrict to positive trace-class operators $X\in\mathcal{C}_{1}^{+}$.
Lemma A.2 (Rank-one calibration).
Let $P$ be a rank-one orthogonal projection with $P\in\mathcal{C}$. Define
$h(\lambda):=\mathcal{E}(\lambda P),\qquad\lambda\in[0,\infty).$
Then:
1. $h$ is well-defined (independent of the choice of rank-one $P$); 2. $h$ is nondecreasing on $0,\infty)$; 3. $h$ is Borel-measurable on $0,\infty)$; 4. $h(0)=0$.
Proof.
1. If $P_{1}$ and $P_{2}$ are rank-one projections, there exists a unitary $U$ with $P_{2}=U^{*}P_{1}U$. By unitary invariance,
$\mathcal{E}(\lambda P_{2})=\mathcal{E}\big(U^{*}(\lambda P_{1})U\big)=\mathcal{E}(\lambda P_{1}),$
so $h(\lambda)$ does not depend on the choice of rank-one $P$.
2. If $0\leq\lambda\leq\mu$, then $\lambda P\leq\mu P$ and by positivity and monotonicity of $\mathcal{E}$ (a consequence of dominated continuity applied to the monotone sequence $\lambda_{n}P\uparrow\mu P$), we have
$h(\lambda)=\mathcal{E}(\lambda P)\leq\mathcal{E}(\mu P)=h(\mu).$
Thus $h$ is nondecreasing.
3. Let $\lambda_{n}\downarrow\lambda$. Then $\lambda_{n}P\downarrow\lambda P$ strongly, dominated by $\lambda_{1}P$. By dominated continuity,
$\mathcal{E}(\lambda_{n}P)\to\mathcal{E}(\lambda P),$
so $h$ is right-continuous with left limits. Any monotone function on $\mathbb{R}$ is Borel-measurable.
4. Finally, $0\cdot P=0$, and by positivity and projector-locality, $\mathcal{E}(0)=0$. Hence $h(0)=0$. ∎
Thus we have a canonical Borel-measurable nondecreasing function $h:[0,\infty)\to[0,\infty)$ determined uniquely by the values of $\mathcal{E}$ on rank-one spectral blocks.
We next extend the finite-dimensional argument to finite-rank and then to general trace-class inputs.
Lemma A.3 (Finite-rank operators).
Let $X\in\mathcal{C}$ be a finite-rank positive operator with spectral decomposition
$X=\sum_{j=1}^{m}\lambda_{j}P_{j},$
where $\lambda_{j}\geq 0$ and the projections $P_{j}$ are pairwise orthogonal and finite-rank. Then
$\mathcal{E}(X)=\operatorname{Tr}\big(h(X)\big)=\sum_{j=1}^{m}h(\lambda_{j})\,\operatorname{Tr}(P_{j}).$
Proof.
Write each $P_{j}$ as an orthogonal sum of rank-one projections:
$P_{j}=\bigoplus_{k=1}^{r_{j}}P_{j,k},$
with $\operatorname{rank}(P_{j})=r_{j}$. By extensivity,
$\mathcal{E}(\lambda_{j}P_{j})=\sum_{k=1}^{r_{j}}\mathcal{E}(\lambda_{j}P_{j,k})=\sum_{k=1}^{r_{j}}h(\lambda_{j})=r_{j}\,h(\lambda_{j})=h(\lambda_{j})\,\operatorname{Tr}(P_{j}),$
since $\operatorname{Tr}(P_{j})=r_{j}$ for finite-rank projections. By projector-locality,
$\mathcal{E}(X)=\sum_{j=1}^{m}\mathcal{E}(\lambda_{j}P_{j})=\sum_{j=1}^{m}h(\lambda_{j})\,\operatorname{Tr}(P_{j}).$
On the other hand, the functional calculus gives
$h(X)=\sum_{j=1}^{m}h(\lambda_{j})P_{j},$
so
$\operatorname{Tr}\big(h(X)\big)=\sum_{j=1}^{m}h(\lambda_{j})\,\operatorname{Tr}(P_{j}).$
This yields $\mathcal{E}(X)=\operatorname{Tr}(h(X))$. ∎
We now pass from finite rank to arbitrary positive trace-class operators.
Lemma A.4 (Trace-class operators).
For every positive $X\in\mathcal{C}_{1}$,
$\mathcal{E}(X)=\operatorname{Tr}\big(h(X)\big).$
Proof.
Let $X\in\mathcal{C}_{1}^{+}$. By the spectral theorem, there exist nonnegative scalars $\{\lambda_{k}\}_{k\geq 1}$ and pairwise orthogonal finite-rank projections $\{P_{k}\}_{k\geq 1}$ such that
$X=\sum_{k=1}^{\infty}\lambda_{k}P_{k},$
where the series converges in trace norm and
$\sum_{k=1}^{\infty}\lambda_{k}\,\operatorname{Tr}(P_{k})<\infty.$
For each $N\geq 1$ set
$X_{N}:=\sum_{k=1}^{N}\lambda_{k}P_{k}.$
Then $X_{N}\in\mathcal{C}$ by stability of $\mathcal{C}$, $0\leq X_{N}\leq X_{N+1}\leq X$, and $X_{N}\uparrow X$ strongly as $N\to\infty$.
By Lemma A.3 we have
$\mathcal{E}(X_{N})=\operatorname{Tr}\big(h(X_{N})\big)=\sum_{k=1}^{N}h(\lambda_{k})\,\operatorname{Tr}(P_{k}).$
Since $X$ is positive, we have $0\leq X\leq\|X\|_{\mathrm{op}}\,I$. Thus
$0\leq X_{N}\leq\|X\|_{\mathrm{op}}\,I$
for all $N$. Set
$Y:=\|X\|_{\mathrm{op}}\,I.$
By Axiom A5, $Y\in\mathcal{C}$ and
$\mathcal{E}(Y)=\|X\|_{\mathrm{op}}\,\mathcal{E}(I)<\infty.$
The sequence $\{X_{N}\}$ is monotone increasing, satisfies $0\leq X_{N}\leq Y$, and converges strongly to $X$. Dominated continuity (Axiom A4) therefore yields
$\mathcal{E}(X_{N})\to\mathcal{E}(X)\qquad\text{as }N\to\infty.$
On the spectral side, $h$ is nondecreasing and $h(0)=0$. The functional calculus gives
$h(X_{N})=\sum_{k=1}^{N}h(\lambda_{k})P_{k},$
so that
$\operatorname{Tr}\big(h(X_{N})\big)=\sum_{k=1}^{N}h(\lambda_{k})\,\operatorname{Tr}(P_{k})=\mathcal{E}(X_{N}).$
Because the sequence $\{\mathcal{E}(X_{N})\}$ is increasing and bounded above by $\mathcal{E}(Y)$, it converges to $\mathcal{E}(X)$. In particular,
$\sum_{k=1}^{\infty}h(\lambda_{k})\,\operatorname{Tr}(P_{k})=\lim_{N\to\infty}\sum_{k=1}^{N}h(\lambda_{k})\,\operatorname{Tr}(P_{k})=\lim_{N\to\infty}\mathcal{E}(X_{N})=\mathcal{E}(X)<\infty.$
Thus $h(X)$ is a positive trace-class operator with spectral decomposition
$h(X)=\sum_{k=1}^{\infty}h(\lambda_{k})P_{k}$
and
$\operatorname{Tr}\big(h(X)\big)=\sum_{k=1}^{\infty}h(\lambda_{k})\,\operatorname{Tr}(P_{k})=\mathcal{E}(X).$
This establishes the trace-form representation on $\mathcal{C}_{1}^{+}$. ∎
Thus on $\mathcal{C}_{1}^{+}$ the map $X\mapsto\mathcal{E}(X)$ coincides with $X\mapsto\operatorname{Tr}\big(h(X)\big)$. For a general self-adjoint trace-class operator $X\in\mathcal{C}_{1}$, we take its positive and negative parts from the spectral decomposition of $X$; then $X_{\pm}\in\mathcal{C}_{1}^{+}$ and $X=X_{+}-X_{-}$. We define
$\mathcal{E}(X):=\mathcal{E}(X_{+})-\mathcal{E}(X_{-}).$
This extension is well defined because the spectral decomposition of a self-adjoint operator is unique, so $X_{\pm}$ are uniquely determined and lie in $\mathcal{C}_{1}^{+}\subset\mathcal{C}$; dominated continuity for signed inputs follows by applying Axiom A4 separately to the positive and negative parts.
A.3. Normalization and uniqueness
We now match the representation of Lemma A.4 with the statement of Theorem 3.1 and record uniqueness up to a scalar factor.
Lemma A.5 (Normalization and scaling).
Let $h$ be as constructed above. Then:
1. For every positive $X\in\mathcal{C}_{1}$,
$\mathcal{E}(X)=\operatorname{Tr}\big(h(X)\big).$
2. If $\tilde{h}$ is another Borel-measurable function and $\tilde{c}_{\mathcal{E}}>0$ satisfy
$\mathcal{E}(X)=\tilde{c}_{\mathcal{E}}\;\operatorname{Tr}\big(\tilde{h}(X)\big)\quad\text{for all }X\in\mathcal{C}_{1}^{+},$
then there exists $a>0$ such that
$\tilde{h}=a\,h\quad\text{and}\quad\tilde{c}_{\mathcal{E}}=1/a.$
Proof.
(1) has already been established in Lemma A.4 with $c_{\mathcal{E}}=1$; we now justify the freedom to rescale.
For (2), fix a rank-one projection $P\in\mathcal{C}$ and $\lambda>0$. Then
$\mathcal{E}(\lambda P)=\operatorname{Tr}\big(h(\lambda P)\big)=h(\lambda)\,\operatorname{Tr}(P),$
and also
$\mathcal{E}(\lambda P)=\tilde{c}_{\mathcal{E}}\;\operatorname{Tr}\big(\tilde{h}(\lambda P)\big)=\tilde{c}_{\mathcal{E}}\;\tilde{h}(\lambda)\,\operatorname{Tr}(P).$
Since $\operatorname{Tr}(P)>0$, we have
$h(\lambda)=\tilde{c}_{\mathcal{E}}\;\tilde{h}(\lambda)$
for all $\lambda\geq 0$. Thus $\tilde{h}=a\,h$ with $a=1/\tilde{c}_{\mathcal{E}}>0$, and $\tilde{c}_{\mathcal{E}}=1/a$. ∎
Setting $c_{\mathcal{E}}:=1$ in Lemma A.4 yields the simplest normalization. The general form of Theorem 3.1, with arbitrary $c_{\mathcal{E}}>0$ and $h$ unique up to a positive scalar factor, is recovered by replacing $h$ with $ah$ and $c_{\mathcal{E}}$ with $1/a$.
A.4. Measure-theoretic formulation (optional viewpoint)
For completeness, we record the scalar-measure formulation of the evaluator, which parallels the classical trace-form representation via Radon-Nikodym. This subsection is not required for the proof of Theorem 3.1 on $\mathcal{C}_{1}$, but clarifies the connection with spectral measures and addresses measure-theoretic concerns.
Fix a self-adjoint operator $D$ with spectral projection-valued measure $E_{D}(\cdot)$.
Definition A.6 (Trace and evaluator measures).
Define the trace measure $\nu_{D}$ and evaluator measure $\mu_{D}$ on the algebra of Borel sets $B\subset\mathbb{R}$ with $E_{D}(B)\in\mathcal{C}$ by
$\nu_{D}(B):=\operatorname{Tr}\big(E_{D}(B)\big),\qquad\mu_{D}(B):=\mathcal{E}\big(E_{D}(B)\big).$
By stability of $\mathcal{C}$, all such spectral projections lie in $\mathcal{C}$, and both $\nu_{D}$ and $\mu_{D}$ take values in $0,\infty]$.
Lemma A.7 ($\sigma$-additivity and absolute continuity).
1. The set function $\mu_{D}$ is countably additive on increasing unions: if $B_{n}\uparrow B$ and all $E_{D}(B_{n}),E_{D}(B)\in\mathcal{C}$, then
$\mu_{D}(B_{n})\to\mu_{D}(B).$
2. $\mu_{D}$ is absolutely continuous with respect to $\nu_{D}$: if $\nu_{D}(B)=0$, then $\mu_{D}(B)=0$.
Proof.
(1) Let $P_{n}:=E_{D}(B_{n})$ and $P:=E_{D}(B)$. Then $P_{n}\uparrow P$ in the strong operator topology, and $0\leq P_{n}\leq I$ for all $n$. By normalization, $\mathcal{E}(I)<\infty$, and by stability, $I\in\mathcal{C}$. Thus dominated continuity (Axiom 4) applied to the monotone sequence $\{P_{n}\}$, dominated by $I$, yields
$\mu_{D}(B_{n})=\mathcal{E}(P_{n})\to\mathcal{E}(P)=\mu_{D}(B).$
(2) If $\nu_{D}(B)=0$, then $\operatorname{Tr}(E_{D}(B))=0$. Since $E_{D}(B)$ is an orthogonal projection, this implies $E_{D}(B)=0$. Hence
$\mu_{D}(B)=\mathcal{E}\big(E_{D}(B)\big)=\mathcal{E}(0)=0.$
Thus $\mu_{D}\ll\nu_{D}$.
In particular, $\mu_{D}$ is a nonnegative scalar measure on the Borel $\sigma$-algebra (via Caratheodory extension from the projection algebra), absolutely continuous with respect to $\nu_{D}$. ∎
Remark A.8 (Radon-Nikodym density and Borel measurability).
Since $D$ has discrete spectrum with finite multiplicities, the trace measure $\nu_{D}$ is $\sigma$-finite: for each integer $m\geq 1$ the bounded interval [m,m][-m,m] contains only finitely many eigenvalues of $D$, so $\nu_{D}([-m,m])<\infty$, and
$\mathbb{R}=\bigcup_{m\geq 1}[-m,m]$
is a countable union of sets of finite $\nu_{D}$-measure. On this $\sigma$-finite measure space, the Radon-Nikodym theorem yields a Borel function $w_{D}:\mathbb{R}\to[0,\infty)$ such that
$\mu_{D}(B)=\int_{B}w_{D}(\lambda)\,d\nu_{D}(\lambda)$
for all Borel sets $B$ with $\nu_{D}(B)<\infty$. This expresses the evaluator on spectral projections as a density $w_{D}$ with respect to the trace measure $\nu_{D}$.
Proof of Theorem 3.1.
For trace-class spectral transforms $X=f(D)\in\mathcal{C}_{1}^{+}$, the preceding sections show that
$\mathcal{E}(X)=\operatorname{Tr}\big(h(X)\big)=\int_{\mathbb{R}}h(f(\lambda))\,d\nu_{D}(\lambda).$
Comparing this with
$\mathcal{E}(f(D))=\int_{\mathbb{R}}f(\lambda)\,w_{D}(\lambda)\,d\nu_{D}(\lambda),$
we see that the scalar density $w_{D}$ encodes the same evaluator data as the global transform $h$, but the latter is intrinsic (independent of the choice of $D$) and directly compatible with the functional calculus on $\mathcal{C}_{1}$.
This concludes the proof of Theorem 3.1 and its measure-theoretic interpretation by combining Lemmas A.1, A.2, A.3, A.4, A.5, and A.7. ∎
Appendix B Counting-function asymptotics and growth stability
In this appendix we collect the counting-function and Tauberian estimates used in the spectral growth taxonomy of Definition 3.4 and Theorem 3.5. Throughout we work with self-adjoint operators with discrete spectrum and finite multiplicities, so that the eigenvalues can be listed in nondecreasing order and counted by a monotone step function.
B.1. Counting functions, asymptotic notation, and regular variation
Let $D$ be a self-adjoint operator on a separable Hilbert space $H$ with discrete spectrum in $0,\infty)$ and finite multiplicities. We write its eigenvalues (repeated according to multiplicity) as
$0\leq\lambda_{1}(D)\leq\lambda_{2}(D)\leq\cdots,\qquad\lambda_{n}(D)\to\infty\text{ as }n\to\infty.$
For brevity in this appendix we abbreviate $\lambda_{n}:=\lambda_{n}(D)$. (Since the spectrum is contained in $0,\infty)$, absolute values in counting conditions are redundant in this appendix.)
Definition B.1 (Counting function).
The counting function of $D$ is
$N_{D}(\Lambda):=\#\{n\geq 1:\lambda_{n}(D)\leq\Lambda\},\qquad\Lambda\geq 0.$ (B.1)
By construction $N_{D}$ is nondecreasing, right-continuous, and takes values in $\mathbb{N}\cup\{0\}$. We assume throughout that $N_{D}(\Lambda)\to\infty$ as $\Lambda\to\infty$.
Definition B.2 (Asymptotic notation).
We use standard asymptotic notation.
1. For functions $g,h:(0,\infty)\to(0,\infty)$ we write
$g(t)\sim h(t)\quad(t\to\infty)$
if
$\lim_{t\to\infty}\frac{g(t)}{h(t)}=1.$
2. We write $g(t)=O(h(t))$ as $t\to\infty$ if there exist constants $C,t_{0}>0$ such that $|g(t)|\leq C\,h(t)$ for all $t\geq t_{0}$.
3. We write $g(t)=o(h(t))$ as $t\to\infty$ if
$\lim_{t\to\infty}\frac{g(t)}{h(t)}=0.$
When useful, we may refine $g(t)\sim h(t)$ to
$g(t)=h(t)\bigl(1+o(1)\bigr)\quad\text{or}\quad|g(t)-h(t)|=o\bigl(h(t)\bigr)$
to make the error behaviour explicit.
We next recall the standard notion of regular variation.
Definition B.3 (Regular variation).
1. A measurable function $L:(0,\infty)\to(0,\infty)$ is slowly varying at infinity if
$\lim_{t\to\infty}\frac{L(ct)}{L(t)}=1\quad\text{for every }c>0.$
2. A measurable function $f:(0,\infty)\to(0,\infty)$ is regularly varying at infinity with index $\rho\in\mathbb{R}$ if
$f(t)=t^{\rho}L(t),$
where $L$ is slowly varying. We then write $f\in\mathrm{RV}_{\rho}$.
In the growth taxonomy, regular variation controls the interaction between the counting function $N_{D}$ and monotone reparametrizations $f(D)$.
For monotone transforms we fix a generalized inverse.
Definition B.4 (Generalized inverse).
Let $f:[0,\infty)\to[0,\infty)$ be nondecreasing and unbounded. Its generalized inverse is the function $f^{-1}:[0,\infty)\to[0,\infty)$ defined by
$f^{-1}(\Lambda):=\sup\{t\geq 0:f(t)\leq\Lambda\},\qquad\Lambda\geq 0,$
with the convention that $\sup\emptyset=0$. Then $f^{-1}$ is nondecreasing and right-continuous, and for each $\Lambda$ the value $f^{-1}(\Lambda)$ is the largest $t$ such that $f(t)\leq\Lambda$. In particular, for all $\lambda\geq 0$,
$\lambda\leq f^{-1}(\Lambda)\quad\Longleftrightarrow\quad f(\lambda)\leq\Lambda.$
This is the convention used in Lemma B.5 to express the counting function of $f(D)$ directly in terms of $N_{D}$ and $f^{-1}$.
B.2. Monotone functional calculus and counting functions
We now justify the counting-function relation used in Lemma 4.1 and the monotone functional-calculus part of Theorem 3.5.
Let $D$ and $N_{D}$ be as above, and let $f:[0,\infty)\to[0,\infty)$ be bounded on compact intervals and nondecreasing, with $f(\Lambda)\to\infty$ as $\Lambda\to\infty$. By the spectral theorem, $f(D)$ is self-adjoint with discrete spectrum $\{f(\lambda_{n})\}_{n\geq 1}$, again listed in nondecreasing order with multiplicity. Define
$N_{f(D)}(\Lambda):=\#\{n\geq 1:f(\lambda_{n})\leq\Lambda\},\qquad\Lambda\geq 0.$
Lemma B.5 (Counting functions under monotone reparametrization).
With $f^{-1}$ as in Definition B.4 we have, for all $\Lambda\geq 0$,
$N_{f(D)}(\Lambda)=N_{D}\bigl(f^{-1}(\Lambda)\bigr).$
In particular, if $N_{D}(\Lambda)\to\infty$ as $\Lambda\to\infty$, then
$N_{f(D)}(\Lambda)\sim N_{D}\bigl(f^{-1}(\Lambda)\bigr)\qquad(\Lambda\to\infty).$
Proof.
Fix $\Lambda\geq 0$. By definition of the counting functions,
$N_{f(D)}(\Lambda)=\#\{n\geq 1:f(\lambda_{n})\leq\Lambda\}.$
On the other side, since the eigenvalues $\{\lambda_{n}\}$ are nondecreasing and $f^{-1}(\Lambda)$ is defined as the supremum of $\{t:f(t)\leq\Lambda\}$, we have
$\lambda_{n}\leq f^{-1}(\Lambda)\quad\Longleftrightarrow\quad f(\lambda_{n})\leq\Lambda.$
Indeed, if $\lambda_{n}\leq f^{-1}(\Lambda)$, then by definition of the supremum there exists a sequence $t_{k}\uparrow f^{-1}(\Lambda)$ with $f(t_{k})\leq\Lambda$ for all $k$; the monotonicity of $f$ gives $f(\lambda_{n})\leq\Lambda$. Conversely, if $f(\lambda_{n})\leq\Lambda$ but $\lambda_{n}>f^{-1}(\Lambda)$, then by right-continuity of $f^{-1}$ there exists $\Lambda^{\prime}<\Lambda$ with $f^{-1}(\Lambda^{\prime})\geq\lambda_{n}$, which contradicts the definition of $f^{-1}(\Lambda^{\prime})$. Thus
$\{n\geq 1:f(\lambda_{n})\leq\Lambda\}=\#\{n\geq 1:\lambda_{n}\leq f^{-1}(\Lambda)\}=N_{D}\bigl(f^{-1}(\Lambda)\bigr),$
proving the claimed identity for all $\Lambda\geq 0$.
If $N_{D}(\Lambda)\to\infty$ as $\Lambda\to\infty$, then the same identity, evaluated at $\Lambda_{k}\to\infty$, shows that the leading-order asymptotic of $N_{D}$ along the reparametrized scale $f^{-1}(\Lambda)$ is inherited by $N_{f(D)}$. In particular,
$N_{f(D)}(\Lambda)\sim N_{D}\bigl(f^{-1}(\Lambda)\bigr)$
in the sense of Definition B.2. ∎
This lemma justifies the counting-function relation used in the monotone functional-calculus part of Theorem 3.5: under the discrete-spectrum hypotheses and the monotonicity of $f$, we may treat $N_{f(D)}(\Lambda)$ and $N_{D}(f^{-1}(\Lambda))$ as identical for asymptotic purposes.
When one prefers to state the result with an error term, Lemma B.5 implies
$N_{f(D)}(\Lambda)=N_{D}(f^{-1}(\Lambda))+O(1)\quad\text{and}\quad|N_{f(D)}(\Lambda)-N_{D}(f^{-1}(\Lambda))|=o\bigl(N_{D}(f^{-1}(\Lambda))\bigr)$
as $\Lambda\to\infty$, in the sense of Definition B.2.
B.3. Tensor-product bounds in the polynomial class
We now record the tensor-product estimates used in the polynomial part of Theorem 3.5. Let $D$ and $E$ be self-adjoint operators with nonnegative, discrete spectra and counting functions $N_{D},N_{E}$.
We assume that there exist exponents $d_{D},d_{E}\geq 0$ such that for every $\varepsilon>0$,
$N_{D}(t)=O\bigl(t^{d_{D}+\varepsilon}\bigr),\qquad N_{E}(t)=O\bigl(t^{d_{E}+\varepsilon}\bigr),\qquad t\to\infty,$
and both counting functions are eventually nondecreasing, as in Convention 2.2.
Under these hypotheses $D$ and $E$ belong to the polynomial growth class $\mathcal{G}_{\mathrm{poly}}$ of Definition 3.4.
The eigenvalues of the tensor product $D\otimes E$ are the pairwise products $a_{m}b_{n}$, where $\{a_{m}\}_{m\geq 1}$ and $\{b_{n}\}_{n\geq 1}$ are the eigenvalues of $D$ and $E$, listed with multiplicity. A standard argument shows that the counting function of $D\otimes E$ can be written as a Stieltjes convolution:
$N_{D\otimes E}(\Lambda)=\sum_{m\geq 1}N_{E}\!\left(\frac{\Lambda}{a_{m}}\right)=\int_{(0,\infty)}N_{E}\!\left(\frac{\Lambda}{t}\right)dN_{D}(t),\qquad\Lambda>0.$ (B.2)
We use this representation to obtain a polynomial bound.
Fix $\delta\in(0,1)$. Define
$F(\Lambda):=\int_{1}^{\Lambda}t^{-(d_{E}+\delta)}\,dN_{D}(t),\qquad\Lambda\geq 1.$ (B.3)
Lemma B.6 (Polynomial convolution bound).
Assume that for some $d_{D}\geq 0$ and $\delta\in(0,1)$,
$N_{D}(t)\leq C_{D}\,t^{d_{D}+\delta}\quad\text{for all }t\geq 1,$
for a constant $C_{D}>0$. Then there exists $C_{\delta}>0$ such that
$F(\Lambda)\leq C_{\delta}\,\Lambda^{\max(d_{D}-d_{E}+\delta,\,0)}\quad\text{for all }\Lambda\geq 1.$
Proof.
By Stieltjes integration by parts,
$F(\Lambda)=N_{D}(\Lambda)\,\Lambda^{-(d_{E}+\delta)}-N_{D}(1)+(d_{E}+\delta)\int_{1}^{\Lambda}N_{D}(t)\,t^{-(d_{E}+\delta+1)}\,dt.$
For the boundary term, the growth assumption gives
$N_{D}(\Lambda)\,\Lambda^{-(d_{E}+\delta)}\leq C_{D}\,\Lambda^{d_{D}+\delta}\,\Lambda^{-(d_{E}+\delta)}=C_{D}\,\Lambda^{d_{D}-d_{E}}.$
For the integral term we again use $N_{D}(t)\leq C_{D}t^{d_{D}+\delta}$. Then
$\int_{1}^{\Lambda}N_{D}(t)\,t^{-(d_{E}+\delta+1)}\,dt\leq C_{D}\int_{1}^{\Lambda}t^{d_{D}+\delta}\,t^{-(d_{E}+\delta+1)}\,dt=C_{D}\int_{1}^{\Lambda}t^{d_{D}-d_{E}-1}\,dt.$
If $d_{D}-d_{E}>0$, the integral is $O(\Lambda^{d_{D}-d_{E}})$; if $d_{D}-d_{E}=0$, it is $O(\log\Lambda)$; and if $d_{D}-d_{E}<0$, it is bounded uniformly in $\Lambda$. In the marginal $\log\Lambda$ case, we may absorb the logarithm into $\Lambda^{\delta}$ for any fixed $\delta\in(0,1)$. Thus in all cases there is a $C^{\prime}_{\delta}>0$ such that
$\int_{1}^{\Lambda}N_{D}(t)\,t^{-(d_{E}+\delta+1)}\,dt\leq C^{\prime}_{\delta}\,\Lambda^{\max(d_{D}-d_{E}+\delta,\,0)}.$
Combining the boundary term, the integral term, and the constant $N_{D}(1)$, and absorbing constants into a single $C_{\delta}$, we obtain
$F(\Lambda)\leq C_{\delta}\,\Lambda^{\max(d_{D}-d_{E}+\delta,\,0)}\quad\text{for all }\Lambda\geq 1,$
as claimed. ∎
Remark B.7 (Tauberian interpretation).
When $N_{D}(t)\sim C_{D}t^{d_{D}}$ with $C_{D}>0$, the above bound and its converse can be seen as a special case of Karamata’s Tauberian theorem for monotone functions; see Bingham-Goldie-Teugels [5]*Theorem 1.7.1. Here we only require the one-sided $O$-bound.
We now translate this bound into a growth estimate for the tensor-product counting function. Suppose in addition that
$N_{E}(t)\leq C_{E}\,t^{d_{E}+\delta}\quad\text{for all }t\geq 1$
for some $d_{E}\geq 0$ and the same $\delta\in(0,1)$.
Corollary B.8 (Tensor-product counting function).
Under the assumptions above there exists $C^{\prime\prime}_{\delta}>0$ such that
$N_{D\otimes E}(\Lambda)\leq C^{\prime\prime}_{\delta}\,\Lambda^{d_{D}+d_{E}+2\delta}\quad\text{for all }\Lambda\geq 1.$
In particular, given any $\varepsilon>0$ we may choose $\delta\in(0,1)$ small enough that $2\delta\leq\varepsilon$, and then
$N_{D\otimes E}(\Lambda)=O\bigl(\Lambda^{d_{D}+d_{E}+\varepsilon}\bigr)\quad\text{as }\Lambda\to\infty.$
Proof.
Using the convolution representation (B.2) and the growth bound for $N_{E}$,
$N_{D\otimes E}(\Lambda)=\int_{(0,\infty)}N_{E}\!\left(\frac{\Lambda}{t}\right)\,dN_{D}(t)\leq\int_{(0,\Lambda]}N_{E}\!\left(\frac{\Lambda}{t}\right)\,dN_{D}(t),$
since $N_{E}(\Lambda/t)=0$ for $t>\Lambda$ if $E$ is nonnegative.
For $1\leq t\leq\Lambda$,
$N_{E}\!\left(\frac{\Lambda}{t}\right)\leq C_{E}\left(\frac{\Lambda}{t}\right)^{d_{E}+\delta}=C_{E}\,\Lambda^{d_{E}+\delta}\,t^{-(d_{E}+\delta)}.$
The finitely many eigenvalues of $D$ in $(0,1)$ contribute only $O(1)$ to the integral, so we have
$N_{D\otimes E}(\Lambda)\leq C_{E}\,\Lambda^{d_{E}+\delta}\int_{1}^{\Lambda}t^{-(d_{E}+\delta)}\,dN_{D}(t)+O(1)=C_{E}\,\Lambda^{d_{E}+\delta}F(\Lambda)+O(1).$
Lemma B.6 gives $F(\Lambda)\leq C_{\delta}\,\Lambda^{\max(d_{D}-d_{E}+\delta,\,0)}$, hence
$N_{D\otimes E}(\Lambda)\leq C_{E}\,C_{\delta}\,\Lambda^{d_{E}+\delta}\,\Lambda^{\max(d_{D}-d_{E}+\delta,\,0)}+O(1).$
If $d_{D}-d_{E}+\delta\leq 0$, the exponent in the leading term is at most $d_{E}+\delta$, and we may absorb the $O(1)$ term into the same power. If $d_{D}-d_{E}+\delta>0$, then
$\Lambda^{d_{E}+\delta}\,\Lambda^{d_{D}-d_{E}+\delta}=\Lambda^{d_{D}+d_{E}+2\delta}.$
In either case there exists $C^{\prime\prime}_{\delta}>0$ such that
$N_{D\otimes E}(\Lambda)\leq C^{\prime\prime}_{\delta}\,\Lambda^{d_{D}+d_{E}+2\delta}\quad\text{for all }\Lambda\geq 1.$
Given $\varepsilon>0$, choose $\delta\in(0,1)$ with $2\delta\leq\varepsilon$. Then
$N_{D\otimes E}(\Lambda)\leq C^{\prime\prime}_{\delta}\,\Lambda^{d_{D}+d_{E}+\varepsilon}$
for all $\Lambda\geq 1$, which is the stated $O$-bound. ∎
Remark B.9 (Connection with Theorem 3.5).
The hypotheses of Corollary B.8 can be restated in the language of Section 3.3 as follows. There exist exponents $d_{D},d_{E}\geq 0$ such that for every $\varepsilon>0$,
$N_{D}(t)=O\bigl(t^{d_{D}+\varepsilon}\bigr),\qquad N_{E}(t)=O\bigl(t^{d_{E}+\varepsilon}\bigr)\quad(t\to\infty),$
and both counting functions $N_{D},N_{E}$ are eventually nondecreasing.
This is exactly the condition that $D,E\in\mathcal{G}_{\mathrm{poly}}$ with polynomial indices $d_{D},d_{E}$ in the sense of Definition 3.4: the growth bounds above are the defining property of the polynomial class, and the eventual monotonicity is assumed whenever counting functions are used. Conversely, if $D,E\in\mathcal{G}_{\mathrm{poly}}$ with indices $d_{D},d_{E}$, then these bounds hold and the assumptions of Corollary B.8 are satisfied.
Applying Corollary B.8 under these hypotheses gives
$N_{D\otimes E}(\Lambda)\leq C^{\prime\prime}_{\delta}\,\Lambda^{d_{D}+d_{E}+2\delta}$
for all sufficiently large $\Lambda$ and any fixed $\delta\in(0,1)$. Rewriting $2\delta$ as $\varepsilon>0$ shows that
$N_{D\otimes E}(\Lambda)=O\bigl(\Lambda^{d_{D}+d_{E}+\varepsilon}\bigr)\quad(\Lambda\to\infty)$
for every $\varepsilon>0$, so $D\otimes E\in\mathcal{G}_{\mathrm{poly}}$. This is precisely the tensor-product stability statement in part (3) of Theorem 3.5.
Appendix C Counterexamples to evaluator axioms
This appendix records simple examples showing that the structural axioms imposed on evaluators in Section 3.1 are close to minimal. In particular, we show that dominated continuity and projector-locality cannot be omitted without admitting functionals that are not of trace form.
C.1. Dropping dominated continuity
We construct an evaluator that satisfies unitary invariance, extensivity on orthogonal sums, projector-locality, and normalization, but fails dominated continuity and does not admit a trace-form representation.
Let $H=\ell^{2}(\mathbb{N})$ with standard orthonormal basis $\{e_{n}\}_{n\geq 1}$. For a bounded real sequence $x=(x_{n})_{n\geq 1}$, write $D_{x}=\operatorname{diag}(x_{1},x_{2},\dots)$ for the associated diagonal operator on $H$. Let $\mathcal{C}_{\mathrm{diag}}$ denote the class of all such diagonal self-adjoint operators.
Fix a nonprincipal ultrafilter $\mathcal{U}$ on $\mathbb{N}$. Define a finitely additive set function $\mu$ on subsets of $\mathbb{N}$ by
$\mu(A):=\begin{cases}1,&A\in\mathcal{U},\\ 0,&A\notin\mathcal{U}.\end{cases}$
Then $\mu$ is finitely additive, takes values in $\{0,1\}$, and satisfies $\mu(\mathbb{N})=1$, but is not countably additive.
For a bounded sequence $x=(x_{n})$, define the “ultrafilter limit”
$\mathcal{L}_{\mathcal{U}}(x):=\lim_{n\to\mathcal{U}}x_{n},$
that is, the unique real number $L$ such that for every $\varepsilon>0$,
$\{n\geq 1:|x_{n}-L|<\varepsilon\}\in\mathcal{U}.$
It is standard that such an $L$ exists for bounded real sequences [10].
We now define an evaluator $\mathcal{E}^{\dagger}$ on $\mathcal{C}_{\mathrm{diag}}$ by
$\mathcal{E}^{\dagger}(D_{x}):=\mathcal{L}_{\mathcal{U}}(x).$
We record its basic properties relative to the axioms of Section 3.1.
  1. 1.
    Unitary invariance.
    On the diagonal class $\mathcal{C}_{\mathrm{diag}}$, we restrict to unitaries that permute the standard basis. For such a permutation unitary $U$, the effect on a diagonal operator $D_{x}$ is to permute the entries of $x$. Since $\mathcal{L}_{\mathcal{U}}$ is invariant under finite permutations of the sequence (ultrafilters on $\mathbb{N}$ ignore finite sets), we have
    $\mathcal{E}^{\dagger}(U^{*}D_{x}U)=\mathcal{L}_{\mathcal{U}}(\text{permutation of }x)=\mathcal{L}_{\mathcal{U}}(x)=\mathcal{E}^{\dagger}(D_{x}).$
  2. 2.
    Extensivity on orthogonal sums.
    Let $H=H_{1}\oplus H_{2}$ with the standard decomposition into the first $m$ coordinates and the remaining ones, and let $D_{x}$ and $D_{y}$ act diagonally on $H_{1}$ and $H_{2}$ with sequences $x=(x_{n})$ and $y=(y_{n})$, respectively. The orthogonal sum $D_{x}\oplus D_{y}$ is unitarily equivalent to a diagonal operator $D_{z}$ with sequence $z$ defined by concatenating $x$ and $y$. Finite concatenation modifies only finitely many terms of the tail of the sequence; the ultrafilter limit satisfies
    $\mathcal{L}_{\mathcal{U}}(z)=\mathcal{L}_{\mathcal{U}}(x)=\mathcal{L}_{\mathcal{U}}(y)$
    whenever the tails of $x$ and $y$ agree in $\mathcal{U}$-measure. If, in addition, we restrict to a sub-class where tails do not interact (for instance, by considering block-diagonal operators with disjoint index sets and defining extensivity on the block algebra), we can arrange additivity on orthogonal sums at the level of the algebra generated by such blocks. In particular, on block-diagonal operators $D_{x}\oplus D_{y}$ with orthogonal supports in the index set, we have
    $\mathcal{E}^{\dagger}(D_{x}\oplus D_{y})=\mathcal{E}^{\dagger}(D_{x})+\mathcal{E}^{\dagger}(D_{y}),$
    with appropriate normalization.
    For the purposes of this example, it suffices to note that $\mathcal{E}^{\dagger}$ can be defined on an algebra of diagonal block operators where finite additivity with respect to block decomposition holds.
  3. 3.
    Projector-locality.
    For a subset $A\subset\mathbb{N}$, let $P_{A}=\operatorname{diag}(\mathbf{1}_{A}(n))$ be the associated spectral projection. Then
    $\mathcal{E}^{\dagger}(P_{A})=\mathcal{L}_{\mathcal{U}}(\mathbf{1}_{A}(n))=\mu(A),$
    since the ultrafilter limit of the indicator sequence coincides with $\mu(A)$. For a finite partition $\{A_{j}\}_{j=1}^{N}$ of $\mathbb{N}$ into disjoint measurable pieces, the corresponding spectral decomposition of the identity is
    $I=\sum_{j=1}^{N}P_{A_{j}}.$
    Finite additivity of $\mu$ implies
    $\mathcal{E}^{\dagger}(I)=\mu(\mathbb{N})=\sum_{j=1}^{N}\mu(A_{j})=\sum_{j=1}^{N}\mathcal{E}^{\dagger}(P_{A_{j}}),$
    and the same holds for finite linear combinations of the $P_{A_{j}}$. Thus $\mathcal{E}^{\dagger}$ is projector-local on the diagonal algebra.
  4. 4.
    Normalization.
    The identity operator corresponds to the constant sequence $x_{n}\equiv 1$. Then
    $\mathcal{E}^{\dagger}(I)=\mathcal{L}_{\mathcal{U}}(1,1,\dots)=1.$
We now exhibit the failure of dominated continuity. Consider the increasing sequence of projections
$P_{k}=\operatorname{diag}(\underbrace{1,\dots,1}_{k\ \text{times}},0,0,\dots),\qquad k\geq 1.$
Then $P_{k}\uparrow I$ strongly as $k\to\infty$, and $0\leq P_{k}\leq I$ for all $k$. By the properties of the ultrafilter,
$\mathcal{E}^{\dagger}(P_{k})=\mathcal{L}_{\mathcal{U}}(\mathbf{1}_{\{1,\dots,k\}}(n))=\mu(\{1,\dots,k\})=0$
for each finite $k$, since $\{1,\dots,k\}\notin\mathcal{U}$. On the other hand,
$\mathcal{E}^{\dagger}(I)=1.$
Thus
$\lim_{k\to\infty}\mathcal{E}^{\dagger}(P_{k})=0\neq 1=\mathcal{E}^{\dagger}(I),$
even though the sequence $P_{k}$ is monotone increasing, dominated by $I$, and the dominating operator has finite evaluation $\mathcal{E}^{\dagger}(I)=1$.
This shows that dominated continuity fails for $\mathcal{E}^{\dagger}$. Moreover, $\mathcal{E}^{\dagger}$ is not of trace form: if there were a Borel-measurable nondecreasing $h$ and $c>0$ such that
$\mathcal{E}^{\dagger}(D_{x})=c\,\operatorname{Tr}(h(D_{x}))$
for all bounded diagonal $D_{x}$, then the left-hand side would depend only on the ultrafilter limit of the sequence $x$, whereas the right-hand side depends on the whole sequence via the (possibly regularized) sum of $h(x_{n})$. In particular, sequences with the same ultrafilter limit but different summation behavior would force a contradiction. Thus dropping dominated continuity allows non-trace-form evaluators that are still unitarily invariant, projector-local, and normalized on the diagonal algebra.
C.2. Dropping projector-locality
We next show that projector-locality is an independent requirement. A natural example is provided by the operator norm.
Let $H$ be a separable Hilbert space, and consider the class of bounded self-adjoint operators on $H$. Define
$\mathcal{E}_{\ast}(X):=\|X\|_{\mathrm{op}}$
for bounded self-adjoint $X$, where $\|\cdot\|_{\mathrm{op}}$ denotes the operator norm.
We note the following:
  1. 1.
    Unitary invariance.
    For any unitary $U$,
    $\mathcal{E}_{\ast}(U^{*}XU)=\|U^{*}XU\|_{\mathrm{op}}=\|X\|_{\mathrm{op}}=\mathcal{E}_{\ast}(X),$
    so $\mathcal{E}_{\ast}$ is unitarily invariant.
  2. 2.
    Normalization.
    $\mathcal{E}_{\ast}(I)=\|I\|_{\mathrm{op}}=1$.
  3. 3.
    Failure of projector-locality.
    Let $\{P_{j}\}_{j=1}^{N}$ be a finite family of pairwise orthogonal projections with $\sum_{j=1}^{N}P_{j}=I$, and fix a scalar $c>0$. Then
    $\mathcal{E}_{\ast}(cI)=\|cI\|_{\mathrm{op}}=c,$
    while for each $j$,
    $\mathcal{E}_{\ast}(cP_{j})=\|cP_{j}\|_{\mathrm{op}}=c.$
    Thus
    $\sum_{j=1}^{N}\mathcal{E}_{\ast}(cP_{j})=Nc\neq c=\mathcal{E}_{\ast}(cI)$
    whenever $N\geq 2$. Hence $\mathcal{E}_{\ast}$ fails projector-locality in the strongest possible way: it does not decompose additively over spectral partitions of the identity.
The functional $\mathcal{E}_{\ast}$ is therefore a unitarily invariant, normalized quantity on bounded self-adjoint operators that is not of trace form and does not satisfy the locality axiom. It illustrates that projector-locality is not implied by unitary invariance and basic boundedness properties; it must be imposed explicitly if one wishes to force a trace-form representation.
C.3. Summary
The examples above show that:
  1. 1.
    Without dominated continuity, it is possible to construct evaluators that are finitely additive on spectral projections, unitarily invariant, projector-local on a natural algebra, and normalized, but which fail to be countably additive in the spectral measure sense and are not representable by a trace form.
  2. 2.
    Without projector-locality, even very natural unitarily invariant functionals such as the operator norm fall outside the trace-form framework and cannot be recovered by integrating a density against the spectral trace measure.
Thus the dominated continuity and projector-locality assumptions in Section 3.1 are not redundant: both are needed, in addition to unitary invariance, extensivity, and normalization, to obtain the rigidity result of Theorem 3.1.